### v2

# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

Abstract
A “Non-perturbative General Approximation Scheme”(NGAS) is proposed which is potentially applicable to arbitrary interacting quantum systems described by a Hamiltonian. The essential feature of this scheme consists of realization of a “mapping” of the original “interacting system” on to an “exactly solvable” model, while preserving the major effects of interaction through the self consistency requirement of equal quantum averages of observables in the two systems. We apply the method to the different cases of the one dimensional anharmonic-interactions (AHI), which includes the case of the quartic-, sextic- and octic-anharmonic oscillators and quartic-, sextic-double well oscillators within the harmonic approximation and demonstrate how this simple approach reproduces, in the leading order (LO), the results to within a few percent, of some of the earlier methods employing rather different assumptions and often with sophisticated numerical analysis. We demonstrate the flexibility of the proposed scheme by carrying out the analysis of the AHI by choosing the infinite square-well potential (ISWP) in one dimension as the input approximation. We extend the formalism to ${\lambda\phi^{4}}$-quantum field theory (in the massive symmetric-phase) to show the equivalence of the present method to the “Gaussian-effective potential” approach. The structure and stability of the Effective Vacuum is demonstrated by identifying the Boguliubov-Valatin transformation from the’ free theory’ to the ’effective theory’. A prominent feature of NGAS is the emergence of a new formulation of perturbation theory designated as: “Mean Field Perturbation Theory (MFPT)”, which is free from power-series expansion in any physical parameter, including the coupling strength. Its application is thereby extended to deal with interactions of arbitrary strength and to compute system properties having non-analytic dependence on the coupling, thus overcoming the primary limitations of the “standard formulation of perturbation theory” (SFPT). We demonstrate Borel-summability of MFPT for the case of the quartic- and sextic-anharmonic oscillators and the quartic double-well oscillator (QDWO) by obtaining uniformly accurate results for the ground state of the above systems for arbitrary physical values of the coupling strength. The energy-spectrum obtained for the quartic double-well oscillator may be of special significance since “renormalon”-free, unambiguous results are achieved in MFPT, in contrast to the well-known failure of SFPT in this case. The general nature and the simplicity of the formulation of NGAS lead us to conjecture that this scheme may be applicable to arbitrary interactions in quantum theory.
Keywords: Non-perturbative general approximation scheme; quantum theory of anharmonic- and the double well oscillators; vacuum structure and stability; perturbation theory for arbitrary interaction strength; Borel summability.
pacs:
11.15.Bt,11.10.Jj,11.25.Db,12.38.Cy,03.65.Ge
I Introduction : Need for Approximation in Quantum Theory (QT)
It is well known that quantum theory (QT) describes [1,2] the observed features of the physical universe at the fundamental level, both at the microscopic and the macro- scales. However, in contrast to the vastness of the range of application of QT, the exact analytic solutions to the quantum- dynamics of interacting systems are rather sparse. Some of these cases admitting exact solution are described in the texts [3]. It is not difficult to count and classify the cases where exact analytic solutions are possible. In fact, it is believed [4,5] that the existence of exact analytic solutions is related to the factorization property of the corresponding quantum-Hamiltonian. In that context, recent applications of super-symmetric quantum mechanics (SUSYQM) [6] have resulted in considerable extension [6] of the range and class of exactly solvable potentials in non-relativistic QT.
In view of the rather limited range of physical phenomena amenable to exact analytic solutions in QT, it has become inevitable to develop approximation methods of to derive meaningful results. This aspect was realized soon after the discovery of QT [7].
Approximation methods in quantum theory have a long history-the prominent ones being as old as QT itself [3,7]. There is a vast amount of literature already accumulated over the years on the various approximation schemes (AS) proposed. Nevertheless, the list is ever growing with newer proposals. This could be indicative of the fact that an “ideal” approximation scheme (in the sense of general applicability to the multitude of interacting systems in QT) is perhaps yet to be realized.
In order to motivate and understand the scope of the investigations presented in this report, it is therefore, necessary to review and survey the various existing approximation schemes (AS) in QT. This is attempted in the next section
II Survey and classification of the existing approximation schemes
The main approximation - schemes in QT which have become the most useful and found mention in text books [3] are the following: (a)Perturbation theory (b)Variational method and (c) Semi-classical(WKBJ) approximation schemes. There are several other approaches, some of which, are variants of the basic schemes. These are discussed at length in ref.[8]. However, for completeness, we recount these briefly in the following:
II.1 Perturbation theory (PT)
Perturbation theory (PT) is among the earliest approximation methods recorded in the literature. In the context of classical physics, the method was applied by Lord-Rayleigh [7] as early as 1873. In QT the application of the method is as old as the theory itself, see, e.g. ref.[7].Since the salient features of this method are well discussed in texts (e.g.[3]) only a brief discussion follows.
The standard formulation of PT (SFPT) involves a power-series expansion of observables in some small parameter. Thus, e.g. if $E(g)$ is an observable of the system, then its dependence on the small-parameter denoted here as ‘$g$’, is through the following power-series:
$E(g)=\mathcal{E}_{0}+\sum\limits_{k=1}^{\infty}{g^{k}\mathcal{E}_{k}\equiv\mathcal{E}_{0}(g)+\Delta{\mathcal{E}(g)}}$ (1)
where the first term is exactly analytically computable. The parameter $g$ is usually chosen to be the coupling-strength of interaction defined in a Lagrangian/Hamiltonian frame work. Thus, e.g. if $H$ is the system-Hamiltonian then it can be written in SFPT as:
$H=H_{s}+gH_{I},$ (2)
where $g$ is the coupling-strength of interaction, $H_{s}$ is the exactly solvable unperturbed Hamiltonian and $gH_{I}$ is the ( usually non-linear) interaction part of the Hamiltonian. This latter interaction-term prevents exact analytic solution for the eigen-spectrum of $H$. For the eigen-value(s) of $H$, the perturbation expansion, eqn.(1) has the standard interpretation: ${E(g)}$ is the energy of the interacting-system, $\mathcal{E}_{0}$ is the same without interaction, which can be calculated exactly and the $\mathcal{E}_{k}$ represents perturbative- correction at the $k$-th order to the energy $\mathcal{E}_{0}$ of the non-interacting theory. These corrections are computable, in principle, for arbitrary order $k$.
It is clear from the definition, eqn.(1) that perturbative corrections become meaningful only when these remain sub-dominant to the leading order contribution which is the first term on the r.h.s. Furthermore, the coupling strength, $g$ is restricted to small values, i.e. $|g|<1$, in order that the perturbation series does not become a priori meaningless. A further practical requirement is that the sequence represents either convergent or asymptotic series in order that the eqn.(1) makes sense if truncation of the same after a finite number of terms is to yield results approximating the actual value of the observable. An important corollary is, therefore, that the  rate of convergence of the series ( or, the rate of decrease of the sub-asymptotic sequence, as the case may be ) is an important consideration bearing upon the practical application of the method.
Merits and limitations of perturbation theory
Perturbation theory continues to be the preferred method of approximation for various practical and theoretical reasons [9].
Merits
Two main reasons for its popularity could be :
(i)Provision for systematic order-by-order improvement of the leading-order result.
Perturbation theory is defined ( cf. eqn.(1)) as providing systematic corrections order-by-order over a leading order result, which is exactly computable without any approximation. This could be regarded as the main merit of this method.
(ii) The universality of the method
Perturbation Theory is universal in the sense that it is applicable to a general class of interacting systems provided that one can identify an exactly solvable unperturbed system and a perturbed part involving some small parameter for a power-series expansion. This encompasses a wide class of problems in QT including many-body theory and quantum-field theory (QFT).
Limitations of SFPT
The main limitations of the SFPT could be the following:
(i) Inability to describe the so-called “non-perturbative” phenomena
The nature of the power-series expansion (see eqn.(1)) defining SFPT implies that the latter fails to describe a very important class of phenomena when the interaction-strength $g$ is large: $|g|>1$, as happens, e.g. in strong-interactions, bound states etc.
For the same reason, processes/ phenomena involving non-analytic-dependence on $g$ at the origin remain inaccessible to SFPT. There is a large variety of physical phenomena, which fall under this category, e.g. tunneling , decay, phase-transitions, critical-phenomena, collective and cooperative phenomena such as super-conductivity, super-fluidity etc.
The inability to describe the above class of the so-called “non-perturbative” phenomena remains perhaps the most severe limitation of the method.
(ii)The instability of the “perturbative vacuum”
The “perturbative vacuum/ground state” is defined to be the one belonging to the non-interacting theory. In most of the physical processes it turns out that the true (interacting) vacuum/ ground-state is not analytically accessible from the perturbative vacuum/ ground state when interaction is turned on [8,10]. Furthermore, the instability of the perturbative ground-state has been demonstrated in several cases by establishing that the ground state of the interacting theory lies much lower in energy than the former.
(iii) Inconsistency with the known analytic properties
The analytic properties of an observable such as the energy, at the origin in the complex-$g$-plane have been independently established [11] for some systems and these are inconsistent with those implicit in the defining eqn.(1).
(iv) Practical Difficulties
In addition to the above limitations of the SFPT on theoretical grounds, there may be practical difficulties arising, for example, from computing corrections beyond the first few orders, which generally involve the (infinite) sum over the intermediate states. This problem has been addressed by several authors and it has been found that the difficulty can be surmounted , in specific examples , by appealing to special techniques, such as the Dalgrano-Lewis method [12] and the use of the hyper-virial theorem [9].
Other limitations of PT are discussed in subsequent Section(s) in respect of specific systems.
We close this subsection noting that the difficulties encountered with the SFPT have led to several alternative approaches some of which are discussed below.
II.2 Variational Methods (VM)
The Variational-Method of Approximation (VMA) is a powerful method in obtaining the approximate eigenvalues and eigenstates of observables in QT. Like perturbation theory (PT), the application of the method has a long history- initial application of VMA being ascribed to Lord Rayleigh [13] and to W.Ritz [14]. The VMA becomes especially suitable when perturbation theory (PT) fails or becomes inapplicable to the considered problem. VMA is also used to test results based upon PT and for analysis of stability of the system under consideration.
The essential ingredient of the method is based upon the variational theorem (Theorem-I), which provides an upper bound for the ground state energy of the system :
$\frac{<~{}\psi~{}|~{}H~{}|\psi~{}>}{<~{}\psi|\psi~{}>}~{}\equiv~{}E~{}[\psi]~{}\geq~{}E_{0}~{},$ (3)
where, ${}|\psi>~{}$ is any arbitrary, normalizable state known as the “trial-state”; H is the Hamiltonian, and ${}E_{0}~{}$ is the ground-state energy of the system. ( In the above equation, we have used the standard Dirac-notation for expectation values and also denoted by $E[\psi]$, the energy- functional). The proof of the theorem is based upon straight forward application of eigenfunction expansion method in QT and can be found in any standard texts [3].
For accuracy of estimation it is necessary to obtain an approximation to the least upper bound by judicious choice of the trial-state. This is provided by the Ritz-method [14], which consists in specifying the trial -state ${}|~{}\psi~{}>~{}$ as a function of one or more free-parameters : $\{\alpha_{i}\}$ and then minimizing the energy functional ${}E[\psi~{}]$ with respect to these parameters :
$\partial{E}[\psi]/\partial\alpha_{i}~{}=~{}0~{};~{}~{}\partial^{2}E[\psi]/\partial\alpha_{i}^{2}~{}>~{}0.$ (4)
Thus, the Ritz-method enables one to obtain a least-upper bound (LUB), for a given choice of the trial-state, which is then the closest approximation to the ground-state energy for that choice. Some other relevant aspects of the VMA are discussed below :
(a) The equality sign in eqn.(3) holds if (and only if) the trial-state, $|~{}\psi~{}>~{}$ coincides with the true ground state, $|~{}\psi_{0}~{}>$ of the system.
(b) An error $\sim~{}O(\epsilon)$ in the choice of the trial-state gives rise to an error $\sim~{}O(\epsilon^{2})$ in the computed value of the energy, which means that the system-energy is more accurately determined by the VMA than the wave function.
Consequently, the over-all accuracy of the method sensitively depends on the choice of the trial-state, consistent with the physical boundary conditions for the system. This calls for considerable ingenuity and insight in the choice of the trial-state.
(c) The generalization of the Rayleigh-Ritz method of the VMA to obtain estimates/approximation for the higher excited states can be achieved by several methods. Some of these methods are outlined below :
(i) A tower of trial states, ${}|~{}\psi_{\alpha}~{}>~{};~{}\alpha~{}=~{}1,~{}2,~{}3......,$ approximating the higher excited states of the system may be constructed with the requirement that these are mutually orthogonal to each other, as well as, orthogonal to the variational-ground-state. The construction can be achieved by any suitable method of orthogonalization, such as the Schmidt-method [15] . Denote by $E_{\alpha}$, the energy-functional for the $\alpha$-th state minimized with respect to its free-parameters and then arranged as an decreasing sequence, i.e.,
$E_{\alpha}~{}\equiv~{}\Big{[}~{}\frac{<\psi_{\alpha}~{}|H|~{}\psi_{\alpha}~{}>}{<\psi_{\alpha}~{}|~{}\psi_{\alpha}~{}>}\Big{]}_{min}~{};~{}\alpha~{}=~{}1,~{}2,~{}3,.........$ (5)
The resulting sequence, ${}\mathcal{E}_{0}~{}<~{}E_{1}~{}<~{}E_{2}~{}<~{}E_{3}~{}.......,$ then represents the variational approximation to the energy of the higher excited states sequentially in the order shown.
The other standard method to deal with the excited states within VMA is known as [16] the “method of linear variational approximation (LVA)”. This consists of the choice of the trial-state as a linear superposition of a suitably chosen set of eigenfunctions consistent with the boundary conditions and otherwise appropriate for the system, as described below :
$|\psi>=\sum_{n}c_{n}~{}|u_{n}>~{};\sum_{n}|u_{n}|^{2}~{}=~{}1;~{}~{}<u_{n}|u_{m}>=\delta_{nm}.$ (6)
The energy functional as given in eqn.(3) computed with this trial-state leads to the following set of equations after minimization with respect to the coefficients $c_{n}$ (see, eqn.(6)) :
$\sum_{n}c_{n}^{*}~{}(H_{nm}~{}-~{}E\delta_{nm})~{}=~{}0~{};~{}~{}m~{}=~{}1,~{}2,~{}3,.....$ (7)
where, $H_{nm}\equiv<u_{n}|H|u_{m}>$. Condition for existence of non-trivial solutions of the above set of equation, is given by the vanishing of the secular determinant :
$det~{}(~{}H_{nm}~{}-~{}E\delta_{nm}~{})~{}=~{}0$ (8)
For practical purpose, the determinant has to be truncated at some finite order ‘$k$’. The solution of eqn.(8) then provides, in general, $k$-roots for the energy $E$, which then correspond to the approximate energies of the first ‘${}k~{}$’ excited - states.
It is obvious that the above method becomes efficient in practice when the truncation-error (due to the chosen finite value of $k$) becomes small and further that the spectrum stabilizes (with increasing value of $k$). Clearly, therefore, intelligent choice of the basis-states, $|~{}u_{n}>$ becomes crucial in this approach.
Merits and limitations of the variational method of approximation
The merits and limitations of the VMA may be as described below:
merits
(i) Non-perturbative nature
The method works for arbitrary coupling-strength of interaction as well as for other non-perturbative phenomena with non-analytic dependence on the coupling parameter. This is perhaps the main advantage of VMA.
(ii) Accuracy
Accuracy can be achieved to great precision provided that the trial-states(s) are appropriately chosen. In particular, the diagonalization of the Hamiltonian in a suitably chosen basis-states , with parameters, which are fixed by variational minimization , can attain high accuracy and stability by inclusion of sufficient number of terms in a recursive manner.
(iii) Universality
The method can be applied in principle to arbitrary Hamiltonian systems.
(iv) Establishing vacuum stability
The stability of the ground state/ vacuum of an interacting system can be established by using the basic theorems of the VMA with suitably chosen trial-state. This method has been used e.g. to establish the instability of the perturbative- vacuum in anharmonic-interactions [17], $\lambda\phi^{4}$ QFT [18], QCD [19] etc.
(v)Fail-safe method for guidance
VMA provides guidance to the true ground state and the energy spectrum of the system when no other method is available/ applicable.
Limitations of VMA
(i)Not improvable order-by-order
In the original formulation, this is perhaps the main limitation - unlike PT it is not amenable to systematic improvement order-by-order.
(ii)Non-uniqueness
There is no systematic and standard method to follow in implementation since choice of the trial-states are mainly arbitrary and mostly left to depend upon the ingenuity and insight in the exploration.
(iii) Practical Difficulty in implementation
Practical difficulties may arise in implementation of the VMA when the trial-state(s) involve several variational parameters.
(iv)Not easily generalized to QFT
For QFT, the preferred method of investigation continues to be SFPT mainly because of the advantage of perturbative realization of the renormalization program in a straight forward manner, in spite of valiant attempts within the VMA [18].
There have been several approaches to overcome the limitations of the basic-VMA. These include:(i) improving the input-trial state(s) by either inclusion of more free parameters or otherwise. A sample list can be found in refs.[20-22] for application to standard examples of interacting systems with varying degree of success. However, as noted earlier, these attempts make the variational-minimization of the energy-functional more intractable. (ii)Tightening the variational estimates done by obtaining both an upper-bound as well as, a lower-bound, see e.g. ref.[22]. Lower bounds are obtained by employing standard inequalities and other techniques. (iii)The method of minimum energy variance (MEV): In stead of the Hamiltonian, $H$ itself, upper bounds can be derived, in principle, for any arbitrary function $f(H)$ by the generalization of the standard Rayleigh-Ritz method. When the chosen function is considered as the variance of the Hamiltonian, the method is known as[23] “method of minimum energy variance (MEV)”. The MEV has been shown [23] to lead to improved estimate of the trial-state as well as, the energy eigen-value. In particular, it has been demonstrated[23] that the estimates go beyond the simple Gaussian approximation [17] .
II.3 Semi-Classical Methods (SCM)
The SCM also goes by the name of the original authors: Jeffery-Wentzel-Kramers-Brillouin (JWKB) approximation. The merits of the method are mainly based upon the following features:
Merits (i) model-independent and universal nature : It can be applied to any arbitrary, smooth potential, which covers a large class of interacting systems. (ii) The method is Non-perturbative: It can be applied for arbitrary value of $g$ and also to cases with non-analytic dependence on the latter such as the tunneling / barrier penetration phenomena. (iii) Semi-classical method : The ”Semi-classical” nature of the approximation arises from the fact that the method is based upon the expansion of the “action” of the system as a series involving powers of $\hbar$. Hence the leading-term being $\hbar$-independent, corresponds to the classical-action of the system, which may be easily computable.
The formalism has been dealt at length in texts [24]. Therefore, it will not be repeated here. We note, however, the main features of the approximation scheme as applied to the discrete bound state problems in QT.
The relevant central formula for the above purpose is the so-called JWKB quantization rule :
[…]
where,${}k^{2}(x)~{}=~{}({2m}/{\hbar^{2}})~{}(E~{}-~{}V(x))~{}$; a ,b  are a set of adjacent “turning points” obtained by solving the equation, $k(x)~{}=~{}0~{}$ and other terms have standard meaning. For those cases where the above integral can be evaluated, eqn.(9) can be expressed as :
$f(E,\lambda,g...)~{}=~{}(n+\frac{1}{2})~{}\pi,$ (10)
where, the l.h.s. of eqn.(10) represents the value of the integral as a function of the energy ($E$) and coupling-strengths ${}(\lambda,g....)~{}$ occuring in the potential $V(x)$. The determination of energy $E$ as a function of the other parameters $(n,\lambda,g...)$ therefore, requires the inversion of eqn.(10).
Limitations: With respect to cases covered when eqn.(10) is applicable, the main limitations could be as described below: (i) Non-invertibility of eqn.(10): For most cases of interest , the inversion of eqn.(10) can not be achieved in closed, analytic form except in a few known cases ( which, also happen to be exactly solvable by standard methods). However, eqn.(10) can be inverted by numerical-methods to the desired order of accuracy. (ii)The complexity of the procedure : The complexity of the method increases for the cases, when the WKBJ-integral can not be obtained in closed form as well as, in the cases involving multiple (closely spaced) turning points [25] and the cases requiring higher order of approximation in the WKBJ-series,(e.g. to deal with rapidly varying potentials) become increasingly difficult to implement.
Nevertheless, several useful information /insight about the energy levels can be obtained in limiting cases of the small-coupling regime, the strong-coupling and/or the large - n limits. We refer to Garg [25] who discusses the issues, successes and lists earlier references.
II.4 Other Methods
These are too numerous to be included in this short review. In the following, we succinctly describe the ones that have direct- bearing on the approach presented in this report.
Combination of Variational and Perturbation techniques
Noting that the VMA and the PT broadly complement each other in terms of advantages and limitations, it is not surprising to find in the literature several attempts to incorporate the essential ingredients of the two methods in a single-framework. We mention some of these in the following discussion.
Several authors have worked using this combined approach, including the following: Halliday and Suranyi [26], W.Caswell [27],J.Killingbeck [28], Hsue and Chern [29], Feynman and Kleinert [30] and collaborators [31], Patnaik [32], Rath [33] and many more. The linear delta-expansion method [34], the Gaussian effective potential (GEP) approach [35], the self-consistent field method [36] and its generalizations [37],methods based upon cluster-expansion [38] , coherent- and squeezed states[39] and several other methods utilize the VMA and the VPA in some form or other. The review of each individual work as mentioned above falls beyond the scope of the present report. However, it may be relevant here to note certain common features of the techniques used by these authors and the consequent achievements attained.
In most of the approaches listed above the basic Hamiltonian of the system is altered by the addition and subtraction of terms involving certain additionalparameters. Perturbation techniques can then be applied with redefinition of the unperturbed Hamiltonian and the perturbation correction. The arbitrary parameters are then fixed order-by-order in the resulting perturbation series, by the variational minimization of energy or by imposing other constraints such as the ”principle of minimum sensitivity”(PMS)[40].The resulting sequence of corrections is often claimed [26,27,32,33] to be convergent. Thus, the problem of convergence of the naive perturbation theory as well as the absence of a built-in mechanism of systematic improvement in the VMA, are overcome to a large extent in these hybrid method, which may be called variation-perturbation method (VPM). Apart from the above common underlying feature, each approach differs in detail, which may be found in the individual reference cited above.
The operator Method of Approximation (OMA)
The OMA has been pioneered by Feranchuk and Komarov [41]. The application of the OMA to the anharmonic interaction and to several other problems has been described in ref.[42] , which also provides a guide to earlier works of the authors.
The basic idea of this method is to formulate the problem in the Fock-space of operators, instead of working with the co-ordinates and momenta. In doing so, an arbitrary parameter $\omega$ is introduced into the theory, which is then fixed by standard variational minimization of the energy-functional order-by-order in a modified PT. Details can be found in ref.[41] . The main limitation of the model appears to be the implementation of the scheme e.g. for non-polynomial interactions, interaction representing analytic functions expressible in infinite series, problems in higher dimensions, quantum field theory etc. In such cases, additional assumptions/methods/skills have to be employed [42].
Approximation methods based on quantum-canonical transformation (QCT)
This is a powerful non-perturbative method, which has been successfully applied in QT, particularly in many-body systems exhibiting collective- and co-operative phenomena, e.g. superconductivity and super-fluidity [43-45]. The method had been primarily expounded by Bogoliubov [46] and hence more familiarly known as the “Bogoliubov-Transformation”. These transformations connect the Hillbert-space of the interacting system to that of the interaction-free case, while preserving the canonical structure of the basic ( equal-time) commutation rules . The method is particularly useful in variational studies by proposing the ansatz ( involving the variational - parameters) for the interacting vacuum state (IVS) generated through QCT. After minimization of the energy-functional, the approximation can be tested to dynamically establish the stable ground state of the system. The formalism has been employed in refs.[47,48] for the QAHO and the DWO problems. In the context of ${}\lambda\phi^{4}~{}$ quantum-field theory, QCT has been employed in refs.[49,50,51]. QCT also provides important insight into the structure of the interacting vacuum state [49,50,52,53].
The limitations [54] of the method appear to be those of variational methods as discussed earlier.
Approximation methods based on super symmetric quantum mechanics
The method of super symmetric quantum mechanics (SUSYQM) has been dealt in texts and several review articles. As a representative text [6], may be consulted, which provides guidance to earlier literature.
This has been successfully applied to improve and extend the scope of known approximation methods, e.g. perturbation theory, variational method and JWKB-approximation scheme. In the context of non-relativistic QT, several exact results follow in SUSYQM valid for its partner potential [6] such as the property of iso-spectrality, level degeneracy and positivity. These exact results provide important constraints in testing various approximation schemes.
SUSY-improved perturbation theory starts from an initial guess of the ground state wave function, which can be based, for example, on a realistic variational ansatz. The super-potential corresponding to this trial-ground state wave function can then be constructed by computing the logarithmic derivative of the latter. The unperturbed Hamiltonian is then chosen to be the one obtained from the super potential and the perturbation correction is taken to be the difference of the original Hamiltonian and the unperturbed one. The development of the RS-perturbation theory then becomes straight forward. For illustration of this method, ref.[6] can be consulted.
The SUSY- based JWKB approximation provides the following modification of the quantization-rule :
[…]
where, W is the super-potential and $E^{(1)}_{n}$ is the n-th energy-level of the $H_{1}$, which is one of the partner-Hamiltonians. Similar expression for the energy levels of the other partner-Hamiltonian, $H_{2}$ is given by,
[…]
$~{}n~{}=~{}0,~{}1,~{}2,~{}3,...$ (12)
It may be seen from the above equations that the exact result on the iso-spectrality / level-degeneracy of the partner potentials, is respected in the SUSY-JWKB quantization condition. Moreover, the formulae are valid for all values of ‘$n$’ rather than for large-$n$ as in the case of the conventional JWKB-formulation. The SUSY-JWKB formula is demonstrated [6] to yield results with improved accuracy for many known cases.
In the context of the QAHO/DWO, SUSYQM- methods have been investigated by several authors. A sample list is provided in refs.[55-65].
Approximation Schemes Based upon the Path-Integral Formulation of QT
The Path-Integral (PI) formulation of Quantum Theory [66-68] has traditionally served as an important basis to formulate and improve upon approximation methods. It may be recalled that the Feynman-diagram technique, which is perhaps the most popular method of computation in perturbation theory, originated from the PI-approach [69].The approximation methods can be formulated either in the real-time formulation [66,67] or as the imaginary time, Euclidean-formulation [70]. The latter is found to be ideally suited for extending the application to quantum-statistics [69-71]. A variational method was proposed in the Euclidean formalism of the PI in ref.[71] to obtain an approximation to the partition function of the QAHO and thereby, to obtain an estimate of the ground state energy. Application to quantum field theory and statistical physics /critical phenomena are extensively treated in the text [70].In the context of the QAHO-problem, the stationary-phase approximation was used [72] in the PI-formulation to infer the inapplicability of the SFPT due to the existence of an essential singularity at the origin of the coupling strength plane. In ref.[73], variational lower-bound as well as, upper bound on energy of the QAHO have been obtained by the use of inequalities for the partition function. Semi-classical approximation of partition functions for one-dimensional potentials using the PI-formalism has been described in ref.[74].
Various approximation schemes including the “variational perturbation theory”(VPT)[75,78], semi-classical methods [70,78], Stationary-phase approximation [71,78], loop-wise expansion [78], large-order estimate of perturbation theory [70,78], variational estimation using inequalities [71,78] and many more methods [78] have been based upon the PI-formulation. In particular,the multi-instanton effects for potentials with degenerate ground states is described in ref.[76]. In the context of scalar field theory, ref.[77] describes non-standard expansion techniques for the generating functional in the PI formulation. Several other applications in field theory and statistical physics can be found in ref.[70,78].
The main advantages are wide applications [70,78] and non-perturbative nature of the approximation. Major limitation of the PI-formulation include the mathematical complexity involved in evaluating multi-dimensional integrals - even the simple, exactly solvable potentials such as the hydrogen atom problem, require rather complex techniques [78] for implementation. Inclusion of spin, formulation for fermions and for constrained systems pose special problems. In many cases, it is therefore, preferable to use the standard operator based formalism for implementing the approximation methods. Besides, closed-form analytic expressions for the PI can be obtained only for a few problems with interaction when the former can be approximated by a Gaussian about a saddle-point/ stationary-phase point of the classical action and fluctuations thereabout.
The self-consistent schemes of approximation
These schemes are popular in the context of many-body systems and otherwise known under the name of “mean-field approximation” and the “Hartree-approximation” including various generalizations of the latter, (for example, the “Hartree-Fock” method,“ Hrtree-Fock-Bogoliubov” method etc). The basic concept of the scheme consists in approximating the given potential by an exactly solvable one-body-potential. If the original Hamiltonian is H and the approximating Hamiltonian is $H_{0}$, then perhaps the simplest way to ensure self-consistency in the approximation could be imposition of the following constraint :
$<n|H|n>~{}=~{}<n|H_{0}|n>~{},$ (13)
where, the states $|n>$ are the eigenstates of $H_{0}$. In practice, the approximating Hamiltonian may be chosen with a-priori unknown, adjustable parameters, which then get determined through the constraint, eqn.(13) and other requirements such as the variational principle and / or further simplifying conditions. The non-linear feedback that usually characterizes the self-consistency condition can then be easily ensured. Thus, the simple-looking equation, eqn.(13) possesses the potential to include interaction effects non-perturbatively and the simultaneous preservation of the non-linearity of the original Hamiltonian. The above method of implementation often yields results which are reasonably accurate even in the leading order of approximation, which can be further improved through standard recursive procedure as in the case of the Hartree-Scheme.
In the context of ${}\lambda\phi^{4}~{}$ quantum field theory and the QAHO problem, such a scheme was considered in ref.[79] where, however, the expectation value (see,eqn.(13)) was restricted to the ground state only, in implementing the self-consistency requirement. This limitation has been overcome in proposing a “generalised Hartree-method” in ref.[52]. The ‘mean-field’ approach has also been tried in ref.[80].
The main limitations of the method appear to be the lack of uniqueness of the initial choice of the approximating Hamiltonian and the rate of convergence of the recursive procedure for subsequent improvement.
This completes our survey of the various approximation methods in QT that are relevant in the context of the present report. It may be noted that we have described only those schemes which have general applicability. There are however several other methods which are system-specific. These latter often lead to greater accuracy but only at the cost of losing universality of application. Some such schemes will be described in later Sections when we consider specific applications.
III In Quest of an “Ideal” Approximation Scheme
As detailed in the previous Section-II , ref.[8-80] constitute only a partial list of the main approximation methods /schemes for interacting quantum systems and the list is ever growing. This implies, among other things, that we are far from achieving an ideal approximation scheme to deal with various cases of interacting quantum systems. In the quest of an ideal approximation scheme, the following requirements could perhaps form at least a subset of the list of the desirable criteria characterizing it:
A List of Desirable Criteria for an Ideal Scheme of Approximation :
(i)General Applicability and the Universality of the Scheme:
This requirement means that the scheme must be capable of application to general quantum-systems and for arbitrary strength of interactions. Thus, e.g. it should be applicable to quantum mechanics, many-body systems and quantum-field theory both in the non-relativistic and relativistic domains (This criterion is in contrast to “system-specific methods” which are applicable only to the specific system under study and often primarily for achieving high level of accuracy). Additionally it is also required that the scheme should be applicable for arbitrary strength of interaction, thus encompassing both the perturbative- and non-perturbative domains as per standard nomenclature.
(ii)Simplicity of Implementation:
The simplicity of formalism and ease of implementation/ computation are obvious desirable features of an ideal AS.
(iii)Achievement of Reasonable Accuracy:
Consistent with the criterion (i) above , reasonable accuracy, i.e. at least to within a few percent of the exact/ numerical results, may be desired for general applications. However, it would be an added virtue if better accuracy can be achieved still retaining the general nature of the scheme.
(iv)Provision for Systematical Improvement in an Ordered Manner:
The scheme may be capable of achieving further improvement in a systematic manner order-by-order over and above a leading order (LO) result and finally
(v)Achieving Rapidity of Convergence of the Perturbative Corrections:
The sequence of higher order corrections in any perturbative framework should either rapidly converge for a convergent sequence or decrease with sufficient rapidity in magnitude, for an asymptotic sequence, in order to be useful for practical implementation.
As stated earlier, it is believed that the above set of criteria may at least form a part of more comprehensive/ exhaustive requirements, if any, on judging a given approximation scheme as an ’ideal’ one. In the next Section we confront the above set of criteria to the extant schemes of approximation in QT.
IV A critique of the various extant schemes of approximation
It is revealed from the above brief, yet hopefully representative , survey (Section-II) that in spite of considerable progress achieved in tackling various problems in interacting quantum systems, an ‘ideal ’ method of approximation based upon the above set of requirements is perhaps yet to be achieved as apparently none of the extant schemes simultaneously satisfy the desired criteria.
In view of the above scenario, there remains a lot of scope to achieve the goal of formulating an ideal scheme based upon the listed criteria. This is then the main motivation for the investigations presented in this report. We elaborate upon this aspect in the following Section.
V Motivation, objectives and the scope of the present report
The present report is devoted to the investigations towards achieving the above ambitious objective, namely, to develop an “ideal” approximation scheme on the stated criteria of general applicability to quantum mechanics and quantum field theory (QFT) and satisfying the criteria of simplicity of formulation, self-consistency; being non-perturbative in nature, yet amenable to systematic improvement in an improved perturbation theory resulting in finally achieving reasonable accuracy.
For simplicity, the scope of the investigations reported here is limited to the quantum physics of the self-interacting Bosonic- systems. In this class, we include the one dimensional quartic anharmonic oscillator (QAHO), quartic double-well oscillator (QDWO), the sextic- anharmonic oscillator (SAHO) and the sextic double-well oscillators (SDWO), the octic-anharmonic oscillator(OAHO) and the $\lambda\phi^{4}$ quantum field theory in physical dimensions and in the massive symmetric phase. These investigations are contained in subsequent Sections and organized as described below:
V.1 Organization of the subsequent sections
We outline the general formulation of the proposed approximation scheme in Section-VI. Hereinafter we refer to this new formulation as “Non-perturbative General Approximation Scheme”(NGAS). In this Section , the non-perturbative nature and the self-consistency of the method, as well as, its general applicability are highlighted. Application to a class of an-harmonic interactions (AHI) in the leading order (LO) is contained in the next Section-VII. In this class, we consider sequentially the cases of the QAHO, the QDWO, the SAHO , the SDWO and the OAHO. The general aspect of the formulation is demonstrated by identical treatment of all these cases within the harmonic approximation. We compute the energy-spectrum and compare the leading order (LO)-results with those obtained from earlier approaches and also with exact numerical results wherever available. In this comparison, it is demonstrated how this simple approach reproduces, in the leading order (LO), the results to within a few percent, of some of the earlier methods employing rather different assumptions and often with sophisticated numerical analysis. In Section-VIII , we demonstrate the flexibility of the present scheme by carrying out the analysis of the AHI by choosing the infinite square-well potential (ISWP) as the input approximation. The stability and flexibility of the scheme is established by reproducing results of comparable accuracy and preservation of the general features of the scheme with this drastically different input.
Next, we extend the formalism to $\lambda\phi^{4}$-quantum field theory in physical dimensions and in the massive symmetric-phase in Section-IX . The renormalization of the bare-parameters of the scheme is carried out and demonstrated to be equivalent to the “Gaussian-effective potential” approach. In particular, the emergence of the non-triviality of the renormalized theory is demonstrated. The study of the properties of the vacuum state of the effective theory in the LO establishes the non-trivial structure of the former. In particular, the equivalence of the present approach to that employing the Bogoliubov-transformations is demonstrated. Further, it is shown that the vacuum of the effective theory corresponds to a condensate of particle-pairs of the free-theory. Moreover,the instability of the perturbative vacuum is established by showing that the vacuum-energy of the effective theory lies much lower than that of the former.
In Section-X we describe the analytic computation of corrections to the LO results in NGAS to arbitrary order in a new formulation of perturbation theory(PT). We refer to this new formulation as the “mean-field perturbation theory (MFPT)” in what follows. In order to compare and contrast the relative merits of the SFPT and MFPT it is deemed necessary to first describe the general features of the two formulations. These are therefore, discussed in Section-XI and Section-XII , respectively for SFPT and MFPT. A common feature in either formulation is the divergence of the perturbation series (PS). Hence it is necessary to elaborate how to make sense of the divergent PS by extracting a finite ”sum” for the former. Hence, we discuss this aspect in some detail also in Section-XI.
The analytic computation of perturbative corrections is achieved by recursive evaluation of the same using as ”tools” , the Feynman-Hellman theorem (FHT) and the hyper-virial theorem (HVT) . In the context of anharmonic interactions and using the standard formulation of perturbation theory(SFPT), these tools were first used by Swenson and Danforth (SD) [81] and it is described succinctly here in Section-XI. The generalization of the same formalism to MFPT is contained in Section- XIII.
In Section-XIV, we describe the application to anharmonic-interactions by computing the perturbation corrections, by extending the application of the SD-method to the case of MFPT via the Feynman-Hellman Theorem (FHT) and the Virial Theorems (VT). The large-order behavior of perturbation-corrections in MFPT establishes the asymptotic nature of the series as expected but demonstrated to be Borel-summable in all cases of the AHI investigated in this report . The Borel-summability of the divergent PS for the case of the QDWO in MFPT is highlighted in Section-XIV as this result is in sharp contrast to the situation in SFPT where Borel summation is frustrated by the “renormalon” -singularity in the Borel-plane. Also contained in this Section-XIV, is the characterization of the leading singularity in the Borel-plane, inferred from the large-order behavior of the perturbation series . In Section-XV the computation of the total perturbative corrections (TPC) to the LO-results are presented and tabulated for the different cases of the AHI by employing two independent methods: (i) by the method of optimal-truncation (MOT) of the asymptotic-series at the term of least magnitude and (ii) by Borel-summation. In the latter case, we evaluate the Borel-Laplace integral by using the technique of conformal mapping for the analytic continuation of the Borel-series beyond its circle of convergence and obtain highly accurate results for the energy of the ground state over a wide range of the coupling strength in each case of the AHI considered herein. In Section-XVI, we conclude with a summary of results presented here and compare and contrast these with those obtained in other investigations in the literature. Possible further applications of NGAS to several other areas are pointed out in Section-XVII.
VI A Non-perturbative General Approximation Scheme (NGAS) in Quantum Theory
As has been highlighted in Section-V, the objective of the current investigations is to develop an “ideal” approximation scheme with prescribed features outlined in Section-III. For achieving the above objective, we proceed [82] as follows:
VI.1 General Formulation for Hamiltonian systems
For simplicity and concreteness, we confine the formulation to one dimensional self -interacting quantum systems with stationary , non-degenerate energy eigenvalues. Consider a generic Hamiltonian $H(g)$ describing such a system where $g$ is the coupling strength of the interaction ( hereinafter we suppress the dependence of $H$ on other variables, for notational simplicity). The first step in the implementation of the scheme consists of choosing an “Approximating Hamiltonian” (AH) denoted by $H_{0}$, which is suitable for the system under investigation and consistent with the basic system-properties such as the symmetries and the given initial- and boundary- conditions for it. Other requirements on the AH are detailed below.
VI.1.1 Prescription of conditions on the approximating Hamiltonian (AH)
The eigenvalue equation to be solved for the system-Hamiltonian is given by:
$H(g)|\psi_{n}(g)\rangle=E_{n}(g)|\psi_{n}(g)\rangle,$
where we choose the stationary-energy, $E_{n}(g)$ of the quantum-system as the observable. The Hamiltonian, $H(g)$ can be split in to an exactly solvable dominant part $H_{0}$ and a sub-dominant perturbation,
$H(g)=H_{0}+H^{\prime}.$
By choice, the AH is defined to be $H_{0}$. Then the implementation of NGAS proceeds by imposing the following conditions on the AH:
(i)The AH, $H_{0}$ is exactly solvable:
This condition requires that eigenvalue equation for $H_{0}$ given by
$H_{0}|\phi_{n}\rangle~{}=~{}{E}_{0}^{n}|{\phi}_{n}\rangle;$ (14)
is exactly solvable, i.e. the eigenstates: $|{\phi}_{n}\rangle$ and the eigenvalues: ${E}_{0}^{n}$ are exactly known. Moreover, the eigenstates are normalized and given by
$\langle\phi_{n}|\phi_{n}\rangle~{}=~{}1$ (15)
We designate this requirement, eqn.(14), as the “condition of exact solvability” (CES).
It is further required that
(ii) $H_{0}$ is chosen to depend upon a set of the free (adjustable) parameters,$\{\alpha_{i}\}$ i.e.
$H_{0}=H_{0}(\{\alpha_{i}\})$
In the next step, we impose the ’condition of equal quantum average’, as described below.
VI.1.2 Condition of equal quantum-average (CEQA)
This condition is defined by the requirement that the AH is constrained to yield the same quantum average (QA) as the original system-Hamiltonian, i.e.
$\langle\phi_{n}|H(g)|\phi_{n}\rangle~{}=~{}\langle\phi_{n}|H_{0}(\{\alpha_{i}\})|\phi_{n}\rangle,$ (16)
for all physical ``$n"$ and ``$g"$ where, $|\phi_{n}\rangle$ is defined by the eigenvalue equation for $H_{0}$, eqn.(14) with ``$n"$ denoting the spectral label. In what follows,eqn.(16) will be referred as the “Constraint of Equal Quantum Average (CEQA)”. The next step is to optimize the approximation as described below:
VI.1.3 Determination of the free parameters in AH by variational-optimization
To determine the parameters { ${}\alpha_{i}~{}$}, we impose CEQA as well as variational optimization of the AH in respect of the still undetermined parameters $(\alpha_{i})$:
$\frac{\partial}{\partial{\alpha_{i}}}\langle H_{0}\rangle=0$ (17)
where, the notation is
$\langle{A}\rangle~{}\equiv~{}\langle\phi_{n}|{A}|\phi_{n}\rangle$ (18)
We refer to this condition, eqn.(17) as the “condition of optimality (CO)”.
This simple procedure, not only completely fixes the free parameters ${\alpha_{i}}$ but also generates the feedback of nonlinearity and $g$-dependence in to the AH. This is elaborated in the next subsection.
VI.1.4 The non-linear feedback mechanism of the approximation
The self-consistency of the procedure is implicit in eqs.(14-18). By this procedure the nonlinear $g$-dependence of the system Hamiltonian, $H(g)$ is effectively fed back in to the approximating Hamiltonian, $H_{0}$ such that
$H_{0}(\{\alpha_{i}\})\rightarrow H_{0}(g,n)$ (19)
Thus the above steps generate a mapping from the original ( non-solvable) Hamiltonian $H(g)$ to an exactly solvable AH, $H_{0}$ while preserving the non-linearity of the former in a self-consistent way.
VI.1.5 Features of the Leading order (LO) results in NGAS
The steps which are outlined above are the essential ingredients of the proposed approximation scheme in the leading order(LO). The exact solution of the eigenvalue equation for $H_{0}(g,n)$ given by $E_{0}^{n}(g)$, are naturally identified as the leading-order (LO) result and these are expected to provide the major contribution for the true eigenvalue, $E_{n}(g)$.
At this point, the following observations regarding the approximation scheme are in order:
(i) It is to be noted that in a restricted form, i.e. when the quantum average in eqn.(16) is restricted to the ground state only, the CEQA as expressed in eqn.(16), corresponds to the Hartree-approximation/ mean field approximation in quantum field theory [36] . In this sense NGAS can be regarded as a “generalized” Hartree-approximation method [37]. (ii) It may be emphasized that even the LO results capture the dominant contribution of the (nonlinear) interaction through the self-consistency procedure of NGAS, even though one always deals with an exactly solvable Hamiltonian $H_{0}$. We consider this as a key feature of the approximation method. (iii) Systematic improvement of the LO-result can be implemented in NGAS through the development of an improved perturbaive framework as follows. Consider the following equation which follows from CEQA, eqn.(16):
$\langle\phi_{n}|H^{\prime}(g,n)|\phi_{n}\rangle=0,$ (20)
where
$H^{\prime}\equiv(H-H_{0}).$ (21)
. Note further that
$\langle\phi_{n}|H^{\prime}(g,n)|\phi_{n}\rangle\ll\langle\phi_{n}|H_{0}(g,n)|\phi_{n}\rangle$ (22)
for arbitrary values of $g$ and $n$, implying that the perturbation contribution remains sub-dominant ( in the sense of QA) irrespective of the coupling strength and excitation levels. It is thus naturally suggested to base a perturbation theory treating $H_{0}$ as the unperturbed Hamiltonian and $H^{\prime}$ as the perturbation. This new perturbation theory is expected to possess features that are remarkably different from those in SFPT. The details are discussed Section-X.
VI.2 Section Summary
In this Section-VI, a non-perturbative general scheme of approximation(NGAS) is formulated which is aimed at retaining the merits but overcoming the difficulties/ limitations of the extant methods/ schemes reviewed in the earlier Section-Ii.
The scheme, NGAS is characterized by the following features/properties:
(a) Universal applicability to a general interacting system in Quantum Theory in the Hamiltonian-formulation irrespective of the strength of interaction.
(b) This scheme, by construction, is shown to be potentially capable of providing accurate approximation for any general interaction of arbitrary strength in quantum theory even in the leading order (LO).
(c) Thus, the scheme is essentially non-perturbative in character in the LO.
(d) Yet, it is indicated in this Section-VI how to improve the LO-results still further through the formulation of an improved perturbation theory in a systematic manner within the scheme while retaining the above features of universality and non-perturbative nature.
(e)The basic input in NGAS is the choice of a suitable approximating-Hamiltonian (AH) ${H_{0}}$ , which is required to be exactly solvable, but which involves certain adjustable (variational) parameters, $\alpha_{i}$. The scheme is implemented by imposing the constraint of equal quantum-average (CEQA) such that the quantum average (QA) of the original Hamiltonian $H$ be equal to that of the AH ${H_{0}}$ with respect to any (arbitrary) eigenstate of the latter. The (variational) minimization of this QA with respect to the undetermined free-parameters completely determines the latter.This is referred as the constraint of Optimality (CO).
(f) Apart from the basic requirements as discussed earlier, there is considerable flexibility in the choice of the AH which may be profitably used to improve accuracy of the results.
(g)The evaluation of the energy eigenvalue of the AH determines the leading order(LO) result in NGAS.
(h) The potential development of an improved perturbation theory for arbitrary strength of interaction is contained in NGAS.
In the following Sections, we implement and illustrate the general approach described above by application to specific bench-mark systems which are traditionally employed as theoretical laboratories to test approximation methods in QT.
VII Application of NGAS in the Leading Order (LO) to Anharmonic Interactions (AHI)
Anharmonic Interactions (AHI) are characterized by potential functions that are polynomial additions to the harmonic-interaction. Examples include: quartic anharmonic and double-well potentials, sextic -, octic- and higher anharmonic potentials ( we confine here to parity-symmetric potentials). These can be represented as:
$V(x)=\sum\limits_{k=1}^{K}a_{k}x^{2k}$ (23)
These interactions are quite basic in QT due to their wide applications (see,e.g.[31]) in diverse areas of Physics, as well as, due to their usefulness in traditionally serving as testing grounds for various approximation schemes(see,e.g.[9]) since exact/accurate numerical results are available(see,e.g [8,9]) for comparison. Besides, extensive literature is accessible for comparison of different methods/schemes of approximation applied to the AHI. These are therefore the primary reasons for application of the NGAS to the AHI.
In the following sections, we describe the application of the NGAS in LO-approximation, to the individual AHI-systems.
VII.1 The Harmonic Approximation
Since the AHI are defined by the modification of the harmonic oscillator potential by addition of further polynomial terms of higher degree, it is natural to choose the approximating Hamiltonian (AH), $H_{0}$ as that for the harmonic oscillator(HO) but possibly with a shift in the field-variable $x$ as well as, in the energy. The frequency, $\omega$ of the HO becomes the other free parameter. The exact expression is provided in the subsections that follow. This choice, then defines the Harmonic Approximation, for the different cases of the AHI.
However, as noted earlier, the NGAS is flexible to the choice of the input AH subject only to the defining properties for the latter. In a later section (Section-VIII), we will be using a different approximation, namely the Infinite- Square-Well-Approximation to illustrate this flexibility.
In the next subsections we consider the different cases of the AHI in turn.
VII.2 The Quartic Anharmonic Oscillator (QAHO)
The quartic-anharmonic oscillator ( QAHO ) is the simplest system  exhibiting  self-interaction. This  system has been extensively investigated leading to a vast amount of literature [8,9]. Its importance arises due to physical applications in areas including field theory, condensed matter physics[83], statistical mechanics [84], non-linear systems [85], classical and quantum chaos [86], inflationary cosmology [87], lattice dynamics, plasma oscillations etc, to cite only a few cases. Besides, the QAHO has also served as a theoretical laboratory  to study convergence of perturbation theory [88], development of non-perturbative approximation methods [89], renormalization [90], vacuum structure [91] and stability analysis [92] etc. This quantum system is the basis in testing various aspects of approximation methods, see e.g. [9-11,17-18,21-23,26-42,47-60].
The Hamiltonian of the system is given by
$H~{}=~{}\frac{1}{2}p^{2}~{}+~{}\frac{1}{2}~{}x^{2}+gx^{4},$ (24)
where ${}g~{}$ is real and positive (We deal with dimensionless variables throughout by using scaling in the standard fashion).
In order to develop the NGAS for the QAHO, we follow the steps outlined in the Section-VI, as described below.
(i) Choice of the Approximating Hamiltonian (AH) and the Solution for its spectrum
To obtain the exact analytic solution of the spectrum of ${}H_{0}~{}$ we proceed as follows. We choose the “harmonic approximation” for the input approximating Hamiltonian (AH) defined by the following equation
$H_{0}~{}=~{}\frac{1}{2}p^{2}~{}+~{}\frac{1}{2}~{}\omega^{2}x^{2}+h_{0}.$ (25)
It may be atonce recognized that the AH given by eqn.(25) corresponds to a “shifted” effective harmonic oscillator where the energy is displaced by an amount: $h_{0}$. Note further that the parameter $\omega$ is restricted by physical requirement, to satisfy ${}\omega>~{}0$.
For the implementation of CEQA, one needs to evalute:
$\langle H\rangle=\frac{1}{2}\langle p^{2}\rangle+\frac{1}{2}\langle x^{2}\rangle+g\langle x^{4}\rangle$ (26)
where now the notation is $\langle A\rangle\equiv\langle\phi_{n}|A|\phi_{n}\rangle$, for an operator A . Using the definition of $|\phi_{n}\rangle$ as the eigenfunction of $H_{0}$, the relevant operator-averages occurring in $\langle H\rangle$ can be calculated [8,17] by standard methods , e.g. by using the formalism of the ladder operators for $H_{0}$. These are given by the following equations:
$\langle x^{2}\rangle=(\xi/\omega)$ (27a)
$~{}\langle p^{2}\rangle=\omega\xi,$ (27b)
$\langle x^{4}\rangle=3(1+4\xi^{2})/8\omega^{2},$ (27c)
$\langle x^{6}\rangle=(5/8)(\xi/\omega^{3})(5+4\xi^{2}),$ (27d)
where ${\xi=(n+1/2)}$. Substituting the QA values from eqs.(27a-d) , the QA of the original Hamiltonian $H$ defined by eqn.(26) can now be evaluated as
$\langle\phi_{n}|H|\phi_{n}\rangle\equiv\langle\phi_{n}|H_{0}|\phi_{n}\rangle=\omega\xi/2+(\xi/2\omega)+(3g/8\omega^{2})(1+4\xi^{2})\equiv E_{0}(g,\omega,\xi).$ (28)
(ii) Determination of the free parameters by CEQA and CO
Carrying out the explicit variational minimization of $\langle H_{0}\rangle$ w.r.t ${}^{\prime}\omega~{}^{\prime}$ i.e $\dfrac{\partial\langle H_{0}\rangle}{\partial\omega}=0$ one obtains the following eqn :
$\omega^{3}-\omega-6gf(\xi)=0$ (29)
from where $\omega$ can be obtained as the real, +ve root. It may be emphasized at this point that this equation has been derived by several authors [93], but starting from widely different considerations. Similarly, by noting that $h_{0}=E_{0}(g,\omega,\xi)-\omega\xi$ , the expression for $h_{0}$ is obtained and given by
$h_{0}=\left(\frac{\xi}{4}\right)\left(\frac{1}{\omega}-\omega\right)$ (30)
where $f(\xi)\equiv\xi+\frac{1}{4\xi}$ and $\xi=(n+\frac{1}{2})$ ; with the spctral index $n$ taking values : $n=0,1,2,3,.....$. We refer eqn.(29) as the “ gap equation (GE)” and eqn.(30) as the “energy-shift”. As expected, these parameters thus acquire the functional dependence on $g$ and $\xi$:
${}~{}~{}~{}~{}~{}\omega\rightarrow\omega(g,\xi)~{}~{}~{}~{}~{}~{}h_{0}\rightarrow h_{0}(g,\xi)$ (31)
(iii) Determination of the energy-spectrum in the leading order(LO)
Solution of the gap-equation (GE) eqn.(29) constitutes the key ingredient in the calculation of the energy spectrum. From eqs.(28) and (31), it then follows that the eigen values of $H_{0}$ also acquire the required ${}g~{},~{}\xi~{}$-dependence and the leading order (LO) result for the energy is given by
$E_{0}(g,\xi)=(\xi/4)\left(3\omega+\frac{1}{\omega}\right)$ (32)
where ‘$\omega$’ is obtained as a solution of the $GE$ for the QAHO given by eqn.(29). It is to be noted that the $GE$, eqn.(29) is in the form of a cubic equation of the type
$x^{3}-3Px-2Q=0;~{}P,Q>0$ (33)
The real solution of this eqn.(33) is given by,
$x=Q^{1/3}[(1+\sqrt{(1-\frac{P^{3}}{Q^{2}})})^{1/3}+(1-\sqrt{(1-\frac{P^{3}}{Q^{2}})})^{1/3}]~{}~{}~{}~{}~{}$ (34)
Then comparing the coefficients of eqs.(33) and (29) the required solution is obtained explicitly as
$\omega=(3gf(\xi))^{1/3}[(1+\sqrt{(1-\rho)})^{1/3}+(1-\sqrt{(1-\rho)})^{1/3}],$ (35)
where, ${}\rho^{-1}~{}=~{}243g^{2}f^{2}(\xi)~{}$. It may be noted that, the solution for ‘${}\omega~{}$’ as given above has the correct limiting behaviour, ${}\omega~{}\longrightarrow~{}1$ for ${}g\longrightarrow 0~{}$ and further that it exhibits the non - analytic dependence on the coupling ’${}g~{}$’ at the orgin characteristic of the non-perturbative nature of the NGAS.
At this point, several features/aspects regarding the leading order (LO) results based on Eqs. (29) - (35) may be noted as follows:
(a) As has been mentioned earlier, equation (29) has been derived independently by several authors [93] starting from widely different considerations.
(b) Moreover, the rigorously established [11] analytic structure of $E(g)$ in the $g$ plane, as well as, the non - analytic dependence on the coupling strength ’ $g$’ , are trivially contained [8,82] already at the LO-level through the solution of eqn. (29), which is explicitly given by eqn.(35).
(c) The dependence of $\omega$ and $h_{0}$ on $\xi$ and $g$ implies that the AH, $H_{0}$ also depends on the same parameters. The physical significance of such dependence is that the eigenfunctions of $H_{0}$, $|\phi_{n}\rangle$ are not mutually orthogonal: $\langle\phi_{m}|\phi_{n}\rangle\neq 0$ for $m\neq n$. (As discussed later, this situation does not affect the development of the perturbation corrections in the standard formulation of perturbation theory (SFPT), which can be generated by a ”wave-function- independent method” , such as the Hyper Virial Theorem (HVT) and the Feynman-Hellman Theorem (FVT)-based formalism employed here ( see, Section-X) .
(d) A further significance of $\omega$ arises from the altered ground state structure [8,50,52] due to interaction, differing non-trivially from the free field ground state. Also, the“dressed ground state” of the interacting system in the LO approximation has lower energy [8,50,52,82] as compared to the “trivial” ground state of the noninteracting-theory for any non-vanishing value of the coupling ’$g$’. This result thus establishes the instability of the free-field ground state in presence of interaction. This aspect has been discussed in detail elsewhere [8,50,52,82].
(e) The $(g,\xi)$-dependence of $\omega$ and $h_{0}$ as determined in the LO, does not get altered later in computing the perturbation corrections at higher orders. We elaborate on this important feature later (Section-X).
(f) Further, as stated earlier, the accuracy of the energy spectrum obtained in the LO, eqn.(3.12) is quite significant- the deviations from the “exact” -results being no more than a few percent [8,50,52,82] over the full-range of $g$ and $n$ This result ensures that the dominant contribution indeed comes from the LO as required in a perturbative-framework. We also display the improvement of results obtained by inclusion of the higher-order corrections to the LO-results, which is discussed in Section-X .
We next apply the method to the case of the quartic-double well oscillator in the following subsection.
VII.3 The Quartic Double Well Oscillator (QDWO)
The QDWO is also an extensively studied system [94] because of its theoretical importance and practical applications. The Hamiltonian of the system is given by:
$H~{}=~{}\frac{1}{2}p^{2}~{}-~{}\frac{1}{2}~{}x^{2}+gx^{4};~{}~{}g~{}>0$ (36)
The crucial  -ve   sign of the $x^{2}$- term generates a quite different physical situation than the case of the QAHO, even in the classical limit. The ‘classical’ potential, ${}V_{c}\equiv-\frac{1}{2}x^{2}~{}+~{}gx^{4}~{}$ exhibits the familiar double-well shape with symmetric minima. These minima are located at positions ±${}\pm\frac{1}{2\sqrt{g}}~{}$ and each with depth ${}\frac{1}{16g}~{}$. As ${}g~{}$ becomes smaller and smaller the depth of the two- wells become deeper and deeper. The actual low lying energy eigen-states of the problem become radically different from the trial wave function offered by the harmonic basis. This fact severely handicaps the convergence of the resultant perturbation theory. In principle, the natural solution would be the simultaneous use of two harmonic basis-centers around the positions of the minima at ±${}\pm\frac{1}{2\sqrt{g}}~{}$. However, the implementation of that idea, although possible, implies the use of nonorthogonal states which is rather cumbersome. The   other  difficulty is that the theory is not defined for ${}g\rightarrow 0~{}$, because the ground state does not exist in that limit due to the non-existence of a lower limit to $V_{c}$. In that sense,  the SHO is not the free-field limit of the QDWO. Therefore, the ’standard formulation of perturbation theory (SFPT)’ is not applicable as such, to this case. The perturbation expansion of the eigenvalues ${}E_{n}(g)~{}$ in powers  of ${}g~{}$ is divergent [95-99] for all ${}g~{}>~{}0~{}$. This fact may be understood qualitatively by noting that the addition of the term ${}gx^{4}~{}$ turns a completely continuous eigenvalue spectrum of ${}p^{2}~{}-~{}x^{2}~{}$ into a completely discrete spectrum bounded from below. A non perturbative treatment is therefore, necessary. The WKB method is one such method which is well suited especially for lower energy eigenvalues.
Moreover, it can be mentioned that the case of the QDWO is not Borel-summable [95-99] in SFPT for any value of the coupling strength. Several modern developments such as the theory of resurgence and trans-series [95-97],distributional-Borel summation [98], generalized Borel-Pade method [99] etc. are basically motivated to surmount this problem. In view of the above scenario, the case of the QDWO assumes special relevance for investigation in NGAS.
However, as shown below, in contrast to the situation in SFPT the NGAS can be successfully applied to the case of the QDWO, even in the LO, using completely analogous procedure as in the case of the QAHO, for a considerable larger range of values of ${}n~{}$ and ${}g~{}$.
To develop the NGAS for QDWO the “ approximating Hamiltonian (AH)”for the system is again chosen in the harmonic approximation as:
$H_{0}=(1/2)p^{2}+(1/2){\omega}^{2}(x-\sigma)^{2}+h_{0},$ (37)
but generalized to take into account, the spontaneous symmetry breaking (SSB), through a nonzero vacuum expectation value for $x$ denoted as $\sigma$. Accordingly, the various average values, analogous to eqs.(27a-c) are now given by
$\langle x^{2}\rangle=\sigma^{2}+(\xi/\omega),~{}~{}~{}~{}~{}~{}~{}\langle p^{2}\rangle=\omega\xi,$ (38)
$\langle x^{3}\rangle=\sigma^{3}+3\sigma(\xi/\omega),$ (39)
$\langle x^{4}\rangle=\sigma^{4}+6\sigma^{2}(\xi/\omega)+3(1+4\xi^{2})/8\omega^{2}.$ (40)
Eqs.(38-40) enable the evaluation of $\langle H\rangle$ in terms of the input parameters. These parameters: $\omega$, $\sigma$ and $h_{0}$ are then determined analogously as in the case of the AHO, in terms of $g$ and $\xi$ However, a distinct feature in the case of the QDWO is the occurrence of “quantum-phase transition(QPT)” [82] governed by a “critical coupling” $g_{c}(\xi)$ given by the expression:
$g_{c}(\xi)=\frac{(2/3)^{3/2}}{3(5\xi-(1/4\xi))}$
such that the “Spontaneously Symmetry Broken)(SSB)” phase is realized with $\sigma\neq 0$ for $g\leq g_{c}(\xi)$, whereas the “symmetry-restored (SR) phase” is obtained with $\sigma=0$ when $g>g_{c}(\xi)$. (Numerically, we have: $g_{c}=0.0907218$ , for the ground state of the DWO).
The transition across $g=g_{c}(\xi)$ being discontinuous, the two phases are governed by distinct expressions for $\omega$ and $E_{0}$, which are not analytically connected. It is, therefore, necessary to consider the two phases separately. However, owing to the rather small value of $g_{c}$ the “SSB phase” exists only over a very limited range of $g$ ,i.e when $0\leq g_{c}\leq 0.0907218$ for the ground state and $g_{c}$ takes still smaller values for the higher excited states, eventually vanishing for large $n$ We therefore, confine here to reporting the results for the “SR phase” only.
In this phase, we have $g>g_{c}(\xi)$; $\sigma=0$; $\omega$ satisfies the gap equation
$\omega^{3}+\omega-6g(\xi+\frac{1}{4\xi})=0$ (41)
The energy levels in this phase, are given by the simple expression:
$E_{0}=(\xi/4)(3\omega-(1/\omega))$ (42)
and $h_{0}=E_{0}-\omega\xi$. In eqn.[42] $\omega$ is the solution of eqn.[41] which can be obtained in analogous manner and is given by
$\omega~{}=~{}(3gf(\xi))^{1/3}[(~{}\sqrt{(1+\rho)}+1)^{1/3}~{}-~{}(~{}\sqrt{(1+\rho)}-~{}1)^{1/3}]$ (43)
where, ${}\rho^{-1}~{}=~{}243g^{2}f^{2}(\xi)~{}$. As in the case of the QAHO, the LO-results $E_{0}$ captures the dominant contribution in reproducing the energy spectrum [8,82] to within a few percent (For the ground state energy, see, TABLE-I and -II in Section-X). In these Tables, we also display the improvement of the LO-results obtained by higher order corrections in the MFPT which is discussed in Section-X .
In the next subsection we consider the case of the sextic-anharmonic oscillator in the NGAS.
VII.4 The Sextic Anharmonic Oscillator (SAHO)
The sextic AHO-system is an example of higher anharmonicity, which is also widely investigated [100]. This system is interesting and important in its own right as it finds application in diverse areas of physics [100]. As a theoretical laboratory, this system together with its double-well counterpart provide perhaps the simplest examples, for which supersymmetric quantum mechanics (SUSYQM) has definite predictions [101] for the energy-levels for specific values of the Hamiltonian-parameters. Hence, various models and approximation methods can be tested against the exact analytic results of SUSYQM. Besides, owing to the higher anharmonicity, the divergence of the SFPT at higher-orders becomes even more severe [102], $E_{n}\sim\Gamma[n(m-1)]$ for $gx^{2m}$ type of AHO. This result endows the sextic-AHO with added importance for testing convergent-approximation methods. It may also be noted that the“Wick-ordering” method of Caswell [27] is not directly applicable to this case since the method generates $x^{4}$ counter-terms which are not present in the original (non-ordered) Hamiltonian. For the above stated reasons of practical applicability and theoretical importance, the sextic-AHO provides a unique testing ground for the NGAS, as described in the following.
The Hamiltonian for the system is given by
$H=\frac{1}{2}p^{2}+\frac{1}{2}x^{2}+gx^{2K},~{}~{}~{}~{}(K=3)$ (44)
Results for the SAHO can be obtained by following analogous procedure as in the case of QAHO along with the following input [82] for $\omega$, $h_{0}$ and $E_{0}$: $\omega$ is given by the real, positive root of the ’gap-equation’
$\omega^{4}-\omega^{2}-(15g/4)(5+4\xi^{2})=0.$
One then obtains [82] the following simple expression for the energy-levels of the sextic AHO in the LO:
$E_{0}=(\xi/3)(2\omega+(1/\omega));~{}~{}~{}~{}\\ h_{0}=(\xi/3)\left(\frac{1}{\omega}-\omega\right).$
Again, the LO-results for the energy are obtained accurately [82] to within a few percent over a wide range in $\xi$ and for arbitrary physical values of $g\geq 0$. Further improvement in accuracy is achieved by application of the MFPT as discussed in Section-X .
VII.5 The Octic- Anharmonic Oscillator (OAHO)
The octic-anharmonic oscillator, like its quartic- and sextic-counterparts, finds applications in modelling molecular physics,lattice-vibrations in solids and in quantum chemistry. Because of the higher anharmonicity, the system also provides the theoretical laboratory for more stringent-tests for non-perturbative approximation schemes in quantum theory since the divergence of the SFPT at higher orders of perturbation, becomes still more severe [102] in this case.
The system is considered here to test the generality, uniform applicability and the reliability of NGAS to the case of higher anharmonicities in the AHI. For this case, the detailed description follows:
The Hamiltonian is given by
$H~{}=~{}\frac{1}{2}p^{2}+\frac{1}{2}x^{2}+gx^{8};~{}g~{}>0.$ (45)
We employ in this case, the “approximating Hamiltonian (AH)” given by,
$H_{0}~{}=~{}\frac{1}{2}p^{2}+\frac{1}{2}\omega^{2}(x-\sigma)^{2}+h_{0}$ (46)
Eqn.[46] corresponds to “shifted” effective Harmonic oscillator. The energy spectrum , as in the earlier cases, is
$E_{n}^{(0)}~{}=~{}\omega~{}\xi+h_{0}~{},$ (47)
where, as before:    $\xi~{}=~{}(n+1/2)~{}$ ;${}n~{}=~{}0,1,2,.....$ It is next required to determine the frequency ‘$\omega$’ and ‘$h_{0}$’. For this purpose, we note that the quantum average of the eqn.(45) is given by:
$<H>~{}=~{}<H_{0}>~{}=~{}\frac{1}{2}<p^{2}>+\frac{1}{2}<x^{2}>+g<x^{8}>$ (48)
where the quantum-averages on the r.h.s. of eqn.(48) can be calculated using the standard properties of creation-/annihilation operators as in earlier cases (see, eqn.(27)). After applying the variational minimisation condition and considering $\sigma=0$ to get “physically acceptable” solution, we are led [82] to the simplified GE, given by:
$\omega^{5}-\omega^{3}-35gh(\xi)=0$ (49)
where,               $h(\xi)=\xi^{3}+(7/2)\xi+(9/16\xi)$.
Solution of the Gap Equation and Determination of the Energy Spectrum
The solution of eqn.(49) determines the frequency ‘$\omega$’ of the “shifted” harmonic oscillator. To obtain the energy levels one substitutes ${}\sigma~{}=~{}0$ and ‘$\omega$’ as the solution of eqn.(49). This leads, after some simplification, to the following simple formula[82]:
$E_{0}=(\frac{\xi}{8})(5\omega+\frac{3}{\omega}).$ (50)
where ‘$\omega$’ is obtained by solving eqn.(49) numerically. We compare the LO- result in NGAS [82] with earlier computations [103] over a wide range of values of ‘$g$’ and `$n^{\prime}$. It can be seen from this comparison [82] that the results obtained in the LO of NGAS are already quite accurate over the full range of the parameters, which demonstrates the generality of the method and uniformity of the approximation with increasing anharmonicity.
As stated earlier, an important aspect of NGAS is its flexibility for accommodating different choices for the input Hamiltonian, AH in order to compare and improve up on the final results for the system-Hamiltonian. We demonstrate this feature in the next Section.
VIII Flexibility of Input in NGAS - The square-well approximation
In the previous Sections, the ’simple harmonic-approximation’ was considered as the natural choice for the input/ approximating Hamiltonian (AH) in NGAS for dealing with different cases of AHI. In this Section, we exploit the freedom of choice of the AH in NGAS to consider the same corresponding to the infinite-square-well (ISW)-potential. One important motivation in selecting the ISW-potential as the input, is the pedagogical aspect, as well as the simplicity of this system. It is well-known that the ISW-potential constitutes one of the simplest systems admitting exact analytical solution. As such, it is included in any standard course of introductory quantum mechanics. The establishment of an approximation connecting the AHO/DWO system to the ISW-case, may therefore be considered as an illustrative example of application of a standard text-book topic to advanced research.
Admittedly, however, the ISW- approximation is perhaps the crudest among possible choices of AH. This is due to the fact that the system described by the ISW potential propagates freely between the infinite walls, in stark contrast to the actual situation for the AHO/DWO. Nevertheless, we choose it on purpose here, in order to test the robustness and tolerance of NGAS to the crudest possible input approximation, which could be yet another motivation.
Moreover, as a by-product of this study, it is possible to obtain an approximation for the celebrated case of the SHO in the limit as the anharmonic- coupling strength is set to zero in the AHO-Hamiltonian, thus gaining further insight into the accuracy of NGAS.
In the next subsection, we demonstrate the method by applying it to the case of the quartic-AHO choosing the ISW-Hamiltonian as the input. The LO-results for the energy spectrum are obtained and compared with the results from other calculations. We also highlight the non-perturbative aspects and the analytic structure of energy as a function of the quartic-coupling.
VIII.1 The Quartic Anharmonic Oscillator in ISW-approximation
The quartic-AHO is defined by the Hamiltonian earlier (see, eqn.(24)) and reproduced below (with a different equation-number for ease of reference):
${H}~{}=~{}{\frac{1}{2}~{}p^{2}}~{}+~{}{\frac{1}{2}~{}x^{2}}~{}+~{}{g{x^{4}}}$ (51)
where, $g>0$
The input-Hamiltonian for the ISW-potential is defined by:
${H^{SW}_{0}~{}}=~{}{\frac{1}{2}~{}p^{2}}~{}+~{}{V_{SW}}$ (52)
where the potential, $V_{SW}$ is given by:
${V_{SW}(x)}~{}=~{}\left\{\begin{array}[]{rc}\infty&\mbox{$|x|\geq a$}\\ h&\mbox{$~{}~{}|x|<~{}a~{}$},\\ \end{array}\right.$ (55)
[…]
$H^{SW}_{0}~{}\phi^{SW}_{n}(x)~{}=~{}\mathcal{E}_{n}~{}\phi^{SW}_{n}(x),$ (56)
is easily solved subject to the boundary-condition:
$\phi^{SW}_{n}(x)~{}~{}=~{}0~{}~{}~{}~{}~{}\textrm{for}~{}~{}|x|\geq{a},$ (57)
which ensures the physical requirement of absolute confinement of the system between the (infinite) potential barriers. For ${}~{}|x|<a,$ the normalized eigenfunctions, vanishing on the boundary of the potential-well, are given by:
$\phi^{SW}_{n}(x)\equiv\phi^{(-)}_{n}(x)~{}=~{}\frac{1}{\sqrt{a}}~{}sin\left(\frac{n\pi x}{2a}\right);~{}~{}\textrm{for}~{}~{}$
$n=2,4,6,8...;$ (58)
and,
$\phi^{SW}_{n}(x)\equiv\phi^{(+)}_{n}(x)~{}=~{}\frac{1}{\sqrt{a}}~{}cos\left(\frac{n\pi x}{2a}\right);~{}~{}\textrm{for}$
$~{}~{}n=1,3,5,7....$ (59)
In the above equations, the (±)plus-or-minus($\pm$)( ± ) super-scripts correspond to the even(odd)-parity solutions. The energy eigenvalues are trivially obtained by solving the Schrödinger equation and given by
${\mathcal{E}_{n}}~{}=~{}{h}~{}+~{}\frac{n^{2}\pi^{2}}{8a^{2}}$ (60)
[…]
The evaluation of $<H>$ using the above definition,eqn.[59]is straight-forward. Noting that:
$\langle H\rangle~{}=~{}\langle\frac{1}{2}~{}p^{2}\rangle~{}+~{}\langle\frac{1}{2}~{}x^{2}\rangle~{}+~{}\langle gx^{4}\rangle,$ (62)
each term in eqn.(60) can be computed easily by exploiting parity-invariance, which forbids parity-changing transitions, i.e. ${}<\phi_{n}^{+}|H|\phi_{n}^{-}>=0=<\phi_{n}^{-}|H|\phi_{n}^{+}>$.
The result is given below:
$\langle H\rangle~{}=~{}\left(\frac{n^{2}\pi^{2}}{8}\right)\left(\frac{1}{a^{2}}\right)~{}+~{}\left(\frac{1}{6}\right)c_{n}a^{2}~{}+ga^{4}\left(\frac{1}{5}-\frac{4c_{n}}{n^{2}\pi^{2}}\right)$ (63)
where,
$c_{n}~{}\equiv~{}1-\left(\frac{6}{n^{2}\pi^{2}}\right)~{}~{}and~{}~{}{n=1,2,3,4,...}.$ (64)
The minimization of the expression for $<H>$ as given above, with respect to $u\equiv~{}\left(1/a^{2}\right)$ leads to the following equation:
$u^{3}-P\left(n\right)u-Q\left(g,n\right)~{}=~{}0,$ (65)
where,
$P\left(n\right)~{}\equiv~{}\left(\frac{4}{3}\right)\left(\frac{c_{n}}{n^{2}\pi^{2}}\right),$ (66)
$Q\left(g,n\right)~{}\equiv~{}\left(\frac{16g}{n^{2}\pi^{2}}\right){\left(\frac{1}{5}-\frac{4c_{n}}{n^{2}\pi^{2}}\right)}.$ (67)
The real, positive root of eqn.(63) is required on physical grounds. This is given by,
$u=\left\{{\left(\frac{8g}{n^{2}\pi^{2}}\right)\left(\frac{1}{5}-\frac{4c_{n}}{n^{2}\pi^{2}}\right)}\right\}^{\frac{1}{3}}\left[\left\{1+\sqrt{1-\rho}\right\}^{\frac{1}{3}}+\left\{1-\sqrt{1-\rho}\right\}^{\frac{1}{3}}\right],$ (68)
where
$\rho~{}\equiv~{}\left(\frac{4}{27}\right)\left(\frac{P^{3}}{Q^{2}}\right).$ (69)
Substitution of eqn.(63) in eqn.(61) and following eqn.(16) and eqn.(17), one obtains the following simple expression for the energy eigen-values in the LO:
$E_{n}^{LO}~{}=~{}\left(\frac{3n^{2}\pi^{2}}{16}\right)u~{}+~{}\left(\frac{c_{n}}{12}\right)\left(\frac{1}{u}\right),$ (70)
where, $u$ is given by eqn.(66). The remaining parameter, `$h^{\prime}$ can then be determined by substitution of eqn.(68) in eqn.(58), given by:
$h~{}=~{}\left(\frac{n^{2}\pi^{2}}{16}\right)u~{}+~{}\left(\frac{c_{n}}{12}\right)\left(\frac{1}{u}\right)$ (71)
At this point, several observations are in order:
(i) note that the free parameters of the input-Hamiltonian,i.e. `$a^{\prime}$ and `$h^{\prime}$ acquire functional dependence on `$g^{\prime}$ and `$n^{\prime}$ through eqn.(66) and eqn.(69). This in turn , implies that the input-Hamiltonian,$H_{0}^{SW}$ (see,eqn.(52) also becomes a function of $g$ and $n$. An obvious consequence is that the eigenfunctions of $H_{0}^{SW}{(g,n)}$ corresponding to different eigenvalues become non-orthogonal as the latter now depends on the excitation-label ’n’,i.e.
[…]
$~{}for~{}m\neq~{}n,$ (72)
(ii) It is seen from eqn.(66) and eqn.(68) that the energy in the LO is a non-analytic function of $g$ at the origin. Hence, the results expressed in these equations are not accessible to ordinary perturbation theory as a power-series expansion in $g$. In this sense, the LO-results are non-perturbative.
(iii) The cube-root singularity at $g=0$ and the branch-point-structure in the $complex~{}g-plane$ as given in eqn.(68) are in conformity with rigorous derivation [11] of the analytic structure of the energy of the quartic-AHO in the coupling strength-plane using sophisticated tools of complex-analysis. It is remarkable that, the correct analytic-structure in $g$ arises here as a simple consequence of the CEQA and the CO of NGAS( see, eqn.(16) and eqn.(17)). (iv) Further, it needs to be emphasized here that the observed features noted above under (i)-(iii) also hold in the analogous case of the simple-harmonic approximation [82] applied to the same system of the QAHO/QDWO as described here in previous Sections. This fact demonstrates that these features are the inherent properties of the scheme, NGAS, being independent of the choice of the input Hamiltonian.
[…]
It is seen from this tabulation that the LO-results are uniformly accurate to within $\sim(2-12)~{}\%$ of the standard results over the full range of $n_{s}$ and $g$. It is further seen that the accuracy of the LO-approximation increases significantly by the inclusion of the 2nd-order correction in IPT.
Table-I
$n_{s}$ $g$ $E_{n}^{LO}$ $Error(LO)$ $E_{n}^{(2)}$ $Exact$ $Error$
[…]
0 0.1 0.6312 12.886 0.5748 0.5591 2.793
1.0 0.9033 12.386 0.8290 0.8038 3.136
10.0 1.6902 12.311 1.5546 1.5049 3.296
100.0 3.5168 12.310 3.2359 3.1314 3.341
1 0.1 1.9636 10.972 1.8058 1.7695 2.055
1.0 3.0366 10.900 2.7999 2.7379 2.267
10.0 5.9051 10.911 5.4479 5.3216 2.374
100.0 12.4162 10.965 11.4561 11.1872 2.403
2 0.1 3.3470 6.640 3.0943 3.1386 1.412
1.0 5.5818 7.771 5.1796 5.1793 0.006
10.0 11.9046 8.151 10.3893 10.3471 0.408
100.0 23.7124 8.242 22.0163 21.9069 0.499
4 0.1 6.3332 1.812 5.8162 6.2203 6.496
1.0 11.2748 2.634 10.3198 10.9636 5.872
10.0 23.1124 2.838 21.1642 22.4088 5.556
100.0 49.2384 3.139 45.0952 47.7072 5.475
10 0.1 16.8320 2.997 18.5203 17.3519 6.733
1.0 32.1164 2.479 35.3936 32.9326 7.471
10.0 67.1872 2.350 73.8641 68.8036 7.355
100.0 143.8104 2.323 157.9952 147.2267 7.312
Table 1: Table for results for the QAHO in the ISW-approximation. For the explanation of entries under different columns, see text
The energy-eigenfunctions in the LO are those as given by eqn.(56) and eqn.(57) but with the width-parameter,`$a^{\prime}$ appearing in these equations now acquire functional dependence on $g$ and $n$ in accordance with eqn.(66).
We discuss the LO-approximation for the SHO in the next subsection.
VIII.2 The simple harmonic-oscillator(SHO) in the ISW-approximation
The exact analytic results for the quantum-SHO are, well known and given in any standard text on quantum mechanics. There is therefore no need for any approximation for this celebrated example. However, these ‘exact’ results provide a further test of the accuracy and efficacy of the ISW-approximation in NGAS. It is purely for this reason that the SHO case is discussed here as a particular case ($g=0$) of the AHO.
The Hamiltonian of the SHO is simply obtained from that of the AHO, eqn.(51) by setting the quartic-coupling $g$ to zero and given by
${H^{SHO}}~{}=~{}{\frac{1}{2}~{}p^{2}}~{}+~{}{\frac{1}{2}~{}x^{2}.}$ (73)
The substitution: $g=0$ in Eqs.(60-63) lead to the following corresponding results for the SHO:
${u=\sqrt{P},}$ (74)
$E_{n}^{LO}|_{SHO}~{}=~{}\left(\frac{n^{2}\pi^{2}}{8}\right)u~{}+~{}\left(\frac{c_{n}}{6}\right)\left(\frac{1}{u}\right).$ (75)
The ‘depth’ parameter for the case of the SHO is given by,
$h_{SHO}~{}=~{}\left(\frac{c_{n}}{6}\right)\left(\frac{1}{u}\right).$ (76)
The LO-approximation for the energy-levels following from Eqs.(72-73) are tabulated in Table-II for typical values of the level-index, ${}n_{s}$ along with the results after inclusion of the 2nd-order correction in IPT (discussed in Subsection-D ). The accuracy of the approximation for both the cases with respect to the exact analytic result are also shown in the same Table. The entries in the 3rd- and the last column represent estimation of errors for the results in the LO and with inclusion of 2nd-order correction respectively, as compared to the ‘exact’ results in the 5th-column.
Table-II
$n_{s}$ $E_{n}^{LO}$ $Error-LO{(\%)}$ $E_{n}^{(2)}$ $Exact$ $Error{(\%)}$
0 0.5678 13.56 0.5091 0.5 1.808
1 1.6703 11.35 1.5239 1.5 1.578
2 2.6272 5.05 2.3888 2.5 4.462
5 5.3952 1.90 5.6456 5.5 2.609
10 9.9508 5.23 10.3945 10.5 1.155
15 14.493 6.49 15.8155 15.5 1.260
Table 2: The LO-results for the SHO in the ISW-approximation at sample-values of ${}n_{s}~{}$. Other entries in the Table are explained in the text
In this context, it may be interesting to obtain an $asymptotic$ estimate for the accuracy of the LO-result given by eqn.(73) as compared to the `$exact^{\prime}$ result. This is given by
$\lim_{n_{s}\to\infty}\left(\dfrac{E_{n}^{LO}}{E_{n}^{exact}}\right)=\left(\dfrac{\pi}{2\sqrt{3}}\right)\simeq~{}0.9069,$ (77)
which corresponds to an error of approximately $9.31\%$ At finite values of $n_{s}$ the errors are of the same order of magnitude.However, the inclusion of the 2nd-order correction in IPT significantly improves the accuracy of approximation.
The next Subsection is devoted to the discussion of the ISW-approximation for the quartic-DWO.
VIII.3 The QDWO
The Hamiltonian for the QDWO is given by (see, eqn.(36))
${H^{DWO}}~{}=~{}{\frac{1}{2}~{}p^{2}}~{}-~{}{\frac{1}{2}~{}x^{2}}~{}+~{}{g{x^{4}};~{}~{}~{}~{}~{}g>0}$ (78)
Apart from the various applications of the DWO as discussed earlier, there is considerable theoretical/mathematical interest in the system. In particular, the instability at $g=0$ ( due to the non-existence of a physical ground state) prevents the application of the naïve perturbation theory. In some versions of modified-perturbation theory it has been established [95] that power series-expansion in $g$ is not even Borel-summable. However, in the context of NGAS, the ISW-approximation is routinely applicable to the DWO-case as well, by merely a change of sign of the $x^{2}$: $x^{2}\rightarrow~{}-x^{2}$ in the corresponding formulae for the AHO. In particular, the equation analogous to eqn.(63) for the DWO now becomes:
$u^{3}~{}+~{}P\left(n\right)u-Q\left(g,n\right)~{}=~{}0,$ (79)
leading to the $physical$ solution (analogous to eqn.(66)) given by
$u=\left\{{\left(\frac{8g}{n^{2}\pi^{2}}\right)\left(\frac{1}{5}-\frac{4c_{n}}{n^{2}\pi^{2}}\right)}\right\}^{\frac{1}{3}}~{}~{}~{}~{}~{}~{}~{}~{}$
[$~{}\left[\left\{\sqrt{\rho+1}+1\right\}^{\frac{1}{3}}-~{}\left\{\sqrt{\rho+1}-1\right\}^{\frac{1}{3}}~{}\right]~{}~{},$ (80)
The results for the energy in LO are therefore given by:
$E_{n}^{LO}|_{DWO}~{}=~{}\left(\frac{3n^{2}\pi^{2}}{16}\right)u~{}-~{}\left(\frac{c_{n}}{12}\right)\left(\frac{1}{u}\right),$ (81)
However, several authors [105] have found it convenient to measure the energy of the DWO from the bottom of the (symmetric) double-well, which means that a term equal to $1/{16g}$ be added to formula, eqn.(79). We denote this quantity as: $E_{n}^{\textrm{ref}}\equiv~{}E_{n}^{LO}+\left(1/{16g}\right)$.
In Table-III, we present the results for $E_{n}^{\textrm{ref}}$ along with results corrected to include 2nd-order perturbation effects in IPT ( see, Subsection-D below) for a range of values of $g$ and $n_{s}$. The relative accuracy of these results with respect to those from earlier computation are also given in this Table. Also shown, are corresponding results from ref.[105] denoted as $Exact$ under column-6.The entries in the last column correspond to (%$({\%})error$ of results under column-5 with respect to those under column-6. The entries under column-3 represent relative accuracy of the LO-results as compared with the $Exact$ results shown under column-6.
Table-III
[…]
0 0.1 0.5049 7.220 0.4726 0.4709 0.350
1.0 0.6422 11.242 0.5967 0.5773 3.358
10.0 1.5468 12.266 1.4246 1.3778 3.398
100.0 3.4480 12.313 3.1732 3.0701 3.359
1 0.1 0.8252 7.476 0.7857 0.7678 2.333
1.0 2.3101 10.898 2.1366 2.0830 2.573
10.0 5.5457 11.009 5.1180 4.9957 2.449
100.0 12.2471 10.993 11.3007 11.0337 2.419
2 0.1 1.7393 6.392 1.6632 1.6348 1.740
1.0 4.6254 8.741 4.2983 4.2536 1.052
10.0 10.7242 8.378 9.9565 9.8947 0.625
100.0 23.4937 8.292 21.8101 21.6947 0.532
4 0.1 3.8678 4.994 3.6471 3.6836 0.991
1.0 9.9139 3.659 9.0975 9.5641 4.878
10.0 22.4582 3.322 20.5659 21.7365 5.385
100.0 48.9325 3.248 44.7980 47.3929 5.475
10 0.1 12.2697 1.292 12.840 12.4303 3.731
1.0 29.7738 2.139 32.5902 30.4248 7.117
10.0 66.0787 2.327 72.6644 67.6167 7.465
100.0 143.2937 2.361 157.7026 146.6738 7.519
[…]
It can be seen from this tabulation that the ISW-approximation uniformly reproduces the standard results to within $\sim{(2-10)~{}\%}$ for the case of LO and to within $\sim{(0.5-7.5)~{}\%}$ with the inclusion of 2nd-order correction in IPT, which is considered in details in the next subsection.
VIII.4 Method for higher order corrections
In the context of NGAS, the ’improved perturbation theory (IPT)’ is the development of the standard Rayleigh-Schrödinger perturbation series (RSPS) with the choice of the $unperturbed~{}Hamiltonian$ as the input-Hamiltonian, $H_{0}$ with the free-parameters in the latter being determined through PEQA and CO (see,eqs.(16-17)). The $perturbation,H^{{}^{\prime}}$ is then defined by eqn.(21). The following properties of the IPT follow from the above assignment which are distinct from those in the conventional perturbation theory in the following aspects:
(i) Using eqn.(20) and eqn.(59) it is seen that the perturbation-correction remains always sub-dominant (by construction), compared to the unperturbed part in the following average-sense:
$|\langle~{}H_{0}\rangle|~{}~{}>>~{}|\langle~{}H^{{}^{\prime}}\rangle|\equiv~{}~{}0.$ (82)
This property may be contrasted with the corresponding situation in the $conventional$ RSPS. Consider e.g. the case of the AHO where the ‘perturbation’ ultimately prevails over the ’unperturbed’ component of the Hamiltonian, no matter however small the quartic-coupling is.
(ii) Secondly, the 1st-order perturbation-correction identically vanishes due to eqn.(20),
(iii) The IPT is not restricted to small values of the coupling strength $g$ since eqn.(80) holds for arbitrary values of the latter.
(iv) There is no small-parameter naturally associated with the perturbation-however, the latter remains small in the average sense as defined by eqn.(80). Therefore, to keep track of order-by-order corrections in IPT, one can adopt the standard trick of introducing an arbitrary, real but finite parameter, $\eta$ through the substitution: $H^{{}^{\prime}}\rightarrow~{}{\eta}H^{{}^{\prime}}$ and set this parameter to unity after the computation.
(v)Property (iv) further implies the ``$universality^{\prime\prime}$ of application of IPT to arbitrary interaction since the perturbation, $H^{{}^{\prime}}$ does not involve any parameter of the original Hamiltonian as the expansion-parameter for the RSPS.
Because of the property (ii) the first non-trivial correction for the $n^{th}$ energy level starts at the 2nd-order and given by the standard expression:
$\Delta{E_{n}^{(2)}}~{}=~{}\sum_{m\neq~{}n}\dfrac{\langle~{}n|H^{{}^{\prime}}|~{}m\rangle\langle~{}m|H^{{}^{\prime}}|~{}n\rangle}{E_{n}^{(0)}-E_{m}^{(0)}}$ (83)
In eqn.(81), we have used the notation:
[…]
where $H^{{}^{\prime}}$ is defined by eqn.(21) and we have displayed the $(n,g)$-dependence of the relevant quantities. As an example, the perturbation-part of the Hamiltonian for the case of the quartic-AHO is given by
${H^{{}^{\prime}}|_{AHO}}~{}=~{}{\frac{1}{2}~{}x^{2}}~{}+~{}{g{x^{4}}}-{h(n,g)},$ (85)
where $h(n,g)$ is given by eqn.(69).
It may be noted that the RSPS can be derived [106] without invoking the properties of the eigenfunctions of the unperturbed Hamiltonian. Therefore, the non-orthogonality properties as expressed in eqn.(70) do not affect the standard formule of the RSPS.
As has been discussed earlier and noted in the Tables-(I-III), the inclusion of the 2nd-order correction to the energy levels defined as: $E_{n}^{(2)}\equiv~{}E_{n}^{LO}~{}+~{}\Delta{E_{n}^{(2)}}$ significantly improves the accuracy of the approximation in all the cases considered here, viz the AHO, DWO and the SHO. (Computation of still higher-order corrections to energy-levels can be carried out by standard techniques in a straight-forward manner as discussed in Section-X).
Finally, we summarize and discuss the main results of this Section as follows.
VIII.5 Section-VIII - Summary
In summary, we have presented a very simple yet accurate approximation in the framework of the NGAS for the quartic anharmonic and the double-well oscillators using the elementary system of the infinite-square-well(ISW) as the approximating Hamiltonian. In the computation of the energy-spectrum, uniform accuracy is achieved to within a few percent of the exact numerical results even in the leading-order, for arbitrary values of the quartic-coupling, $g$ and level index,$n_{s}$ for all the above systems. This situation may be contrasted with the results obtained using text-book methods of approximations, e.g. the naive perturbation method, the variational calculations, the WKB-method etc. Besides, the formalism naturally reproduces the correct analytic-structure of the energy in the complex-$g$ plane otherwise established through rigorous mathematical analysis. To systematically improve the leading-order results further, an improved perturbation theory is formulated in NGAS which is not restricted to small-coupling-expansion and shown to yield further significant improvement in accuracy with the inclusion of the 2nd-order correction only. The results and the method are directly relevant in a pedagogic-context owing to the extreme simplicity of the scheme (NGAS) as well as, the input ISW-approximation. The method can be extended in a straight forward manner to deal with the cases of higher anharmonicity.
IX Application of the NGAS in the Leading Order to $g\phi^{4}$ QFT in the Massive Symmetric Phase
In view of the successful application of the NGAS to the AHO’s and DWO’s as described in earlier sections, it is but natural to extend the formalism to ${}g\phi^{4}~{}$ quantum field-theory and test the consequences of the scheme. It may be worth while pointing out here that the ${}g\phi^{4}~{}$- field theory in physical dimensions ( and in lower dimensions ) is an important physical system which finds crucial applications in diverse areas of physics, e.g., the standard model of particle-physics (Higgs-mechanism) , cosmology [87], condensed matter physics , phase-transitions and critical phenomena [70] etc. Besides, this theory provides the simplest theoretical laboratory for testing the various approximation schemes in quantum field-theory (QFT). Hence it becomes imperative to test the current approximation scheme NGAS by applying the same to $g\phi^{4}$- QFT.
We, therefore take up this task in the following subsections employing the harmonic approximation for the input/ approximating Hamiltonian (AH):
IX.1 The Harmonic approximation
The Lagrangian for the system is given by
${\cal L}=\frac{1}{2}(\partial_{\mu}\phi)(\partial^{\mu}\phi)-\frac{1}{2}m^{2}\phi^{2}-g\phi^{4}~{},$ (86)
where ${}m^{2}~{}>~{}0~{}$. This Lagrangian therefore describes the massive, symmetric version of the theory. The Hamiltonian density derived from the above Lagrangian is given by
${\cal H}~{}=~{}\frac{1}{2}(m^{2}\phi^{2}+\phi_{t}^{2}+\phi_{\alpha}^{2})+g\phi^{4}$ (87)
where we have defined:
$\phi_{t}\equiv~{}\partial\phi(\vec{x},t)/\partial t~{}and~{}\phi_{\alpha}\equiv\partial\phi(\vec{x},t)/\partial x^{\alpha}~{}.$
IX.2 The approximating Hamiltonian
To formulate the NGAS for the above theory we choose an approximating Hamiltonian (AH) denoted by ${}{\cal H}_{0}~{}$ such that ideally the defining constraints as given by eqs.(16-17) are satisfied. However, unlike the case in quantum mechanics, the PEQA (eqn.(16)) is hard to implement in QFT as this would involve multi-particle-states in evaluating the quantum-averages (QA). Hence, on grounds of simplicity, we relax the condition, eqn.(16) by restricting the QA to be evaluated in the “ few-particle” states only. To be more specific, we proceed as follows.
As before, the “ Approximating Hamiltonian” (AH) ${}{\cal H}_{0}~{}$ is given by the following expression :
${\cal H}_{0}~{}=~{}\frac{1}{2}M^{2}\xi^{2}+\frac{1}{2}\xi_{\alpha}^{2}+\frac{1}{2}\xi_{t}^{2}+h_{0},$ (88)
where,
$\xi(\vec{x},t)\equiv\phi(\vec{x},t)-\sigma~{},$ (89)
In addition, ${}\xi_{t}\equiv\partial\xi/\partial t~{};~{}\xi_{\alpha}\equiv\partial\xi/\partial x^{\alpha}$ etc. Eqn.(86) is at once identified to be the Hamiltonian density of the hermitian scalar-field ${}\xi(\vec{x},t)~{}$. The ’free parameters’ introduced in the AH are: $M$, $\sigma$ and $h_{0}$, which are to be determined as per the prescribed procedure by applying eqn.(16-17). For that purpose, it is first necessary to obtain the spectrum of ${}H_{0}~{}$ which is done as follows.
The diagonalization of the AH given by eqn.(86) is straightforward by using the Fourier expansion in terms of creation- and annihilation operators :
$\xi(\vec{x},t)~{}=~{}\phi(\vec{x},t)-\sigma~{}=~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}[b(\vec{k})e^{-ikx}+b^{\dagger}(\vec{k})e^{ikx}]~{},$ (90)
where
$\Omega_{k}(M)~{}\equiv~{}2(2\pi)^{3}\sqrt{|\underline{k}^{2}|+M^{2}}~{}\equiv~{}2~{}(2\pi)^{3}~{}\omega_{k}(M)~{},$ (91)
[…]
$[b(\vec{k}),b^{\dagger}(\vec{q})]~{}=~{}\Omega_{k}(M)\delta^{3}(\vec{k}-\vec{q}),$ (92)
which is a consequence of the ETCR between the ‘field’ ${}\phi(\vec{x},t)~{}$ and its canonical conjugate momentum : ${}\pi(\vec{x},t)~{}\equiv~{}\partial{\cal L}/\partial\dot{\phi}~{}$, given by :
$[~{}\phi(\vec{x},t),~{}\pi(\vec{y},t)~{}]~{}=~{}i~{}\delta^{3}~{}(\vec{x}-\vec{y}).$ (93)
The energy of the system described by ${}H_{0}~{}$ is obtained by standard methods and given by :
$H_{0}\equiv\int d^{3}\vec{x}{\cal H}_{0}(\vec{x},t)=\frac{1}{2}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}[b(\vec{k})b^{\dagger}(\vec{k})+b^{\dagger}(\vec{k})b(\vec{k})]+\int d^{3}\vec{x}~{}h_{0}$ (94)
The spectrum of the states are analogously obtained and denoted by :${}|vac>,|\vec{p}>,~{}|\vec{p_{1}},~{}\vec{p_{2}}>,...$ etc where the effective vacuum state $|vac~{}>$ is defined by
$b(\vec{k})~{}|vac~{}>~{}=~{}0~{},$ (95)
and the multi particle-states are generated by multiple application of the creation-operator ${}b^{\dagger}(\vec{p})~{}$ on ${}|~{}vac~{}>~{}$:
$b^{\dagger}(\vec{p})~{}|~{}vac~{}>~{}=~{}|~{}\vec{p}~{}>$ (96)
$\frac{b^{\dagger}(\vec{p_{1}})~{}b^{\dagger}(\vec{p_{2}})}{\sqrt{2!}}~{}=~{}|~{}\vec{p_{1}},~{}\vec{p_{2}}~{}>,~{}etc.$ (97)
Let us note that,
    $H_{0}|vac>~{}=~{}E_{0}~{}|~{}vac~{}>~{};H_{0}|\vec{p}>~{}=~{}E_{1}(\vec{p})|\vec{p}>,....~{}$ etc, where ${}E_{0},~{}E_{1}~{}$etc correspond to the energy of the corresponding states.
The implementation of PEQA require the evaluation of the QA of monomials of the field ${}\phi(\vec{x},t)~{}$ such as : $<vac|~{}\phi^{n}(\vec{x},t)~{}|vac>~{}\equiv~{}<\phi^{n}(\vec{x},t)~{}>,$ $<\vec{p}|~{}\phi^{n}(\vec{x},t)~{}|\vec{p}>~{}$etc. We first turn to the evaluation of ${}<\phi^{n}(\vec{x},t)>~{}$. This is readily done using ${}{translational-invariance}~{}$ of the vacuum-state :${}|vac>~{}$ and the ETCR,  as given by eqn.(4.7). Some useful results thus obtained are given below :
$<\phi(\vec{x},t)>~{}=~{}\sigma~{},$ (98)
$<\phi^{2}(\vec{x},t)>~{}=~{}\sigma^{2}~{}+~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}~{}\equiv~{}\sigma^{2}~{}+~{}I_{0}~{},$ (99)
$<\phi^{4}(\vec{x},t)>~{}=~{}\sigma^{4}~{}+~{}6\sigma^{2}I_{0}~{}+~{}3I_{0}^{2}~{},~{}etc.$ (100)
Similarly,
$<\phi_{\alpha}^{2}(\vec{x},t)>~{}=~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}|\vec{k}|^{2},$ (101)
$<\phi_{t}^{2}(\vec{x},t)>~{}=~{}\int\frac{d^{3}\vec{k}}{\Omega_{k}(M)}~{}\omega_{k}^{2}(M)$ (102)
Using the above results and the PEQA, one obtains the following equation :
$M^{2}(g,\sigma)\equiv m^{2}+12g\sigma^{2}+12gI_{0}(M^{2})~{},$ (103)
i.e.,
$M^{2}(g,\sigma)=m^{2}+12g\sigma^{2}+6g\int\frac{d^{3}(\vec{k})}{(2\pi)^{3}\sqrt{|\vec{k}|^{2}+M^{2}}}~{},$ (104)
Eqn.(102) can be interpreted as the generation of the ‘mass-gap’(i.e. shift in the bare-mass) due to interaction. In analogy with the terminology used earlier, we refer eqs.(101-102) as the “gap-equation(GE)” of the theory. This equation plays crucial role in the subsequent discussions.
The physical consequences of the theory in the leading-order (LO), such as the spectrum, renormalization, stability properties and the structure of the effective vacuum are discussed in the following subsections.
IX.3 The Effective Potential
The effective-potential (EP) can be defined as :
$U_{0}(\sigma)~{}\equiv~{}<vac|~{}{\cal H}_{0}~{}|vac>~{},$ (105)
such that
$\sigma~{}\equiv~{}<vac|~{}\phi(\vec{x},t)~{}|vac>.$ (106)
[…]
$<~{}{\cal H}_{0}~{}>~{}\equiv~{}<~{}{\cal H}~{}>~{}=~{}\frac{1}{2}m^{2}<\phi^{2}>~{}+~{}\frac{1}{2}<\phi_{t}^{2}>~{}+~{}\frac{1}{2}<\phi_{\alpha}^{2}>~{}+~{}g<~{}\phi^{4}~{}>,$ (107)
which works out to be :
$<~{}{\cal H}_{0}~{}>~{}=~{}\frac{1}{2}m^{2}(\sigma^{2}~{}+~{}I_{0})~{}+~{}\frac{1}{2}\int\frac{d^{3}(\vec{k})}{\Omega_{k}(M)}~{}(~{}\omega^{2}_{k}(M^{2})~{}+~{}|\vec{k}|^{2}~{})$
$+~{}g(\sigma^{4}+6\sigma^{2}I_{0}+3I_{0}^{2}).~{}$ (108)
This can be rewritten as, ( using ${}\omega^{2}_{k}(M^{2})~{}\equiv~{}|\vec{k}|^{2}~{}+~{}M^{2}$)
$<~{}{\cal H}_{0}~{}>~{}=~{}I_{1}~{}-~{}\frac{1}{2}I_{0}~{}(~{}m^{2}~{}+~{}12g\sigma^{2}~{}+~{}12gI_{0}~{})~{}~{}~{}~{}$
$+\frac{1}{2}m^{2}~{}~{}(~{}\sigma^{2}~{}+~{}I_{0}~{})~{}+~{}~{}~{}~{}~{}$
$g~{}(~{}\sigma^{4}~{}+~{}6\sigma^{2}I_{0}~{}+~{}3I_{0}^{2}~{}),~{}~{}~{}~{}~{}~{}~{}$ (109)
where, we have defined :
$I_{n}(x)~{}\equiv~{}\int~{}\frac{d^{3}\vec{k}}{\Omega_{k}(x)}~{}[~{}\omega_{k}^{2}(x)~{}]^{n},~{}~{}~{}~{}~{}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}n~{}=~{}0,~{}\pm~{}1,~{}\pm~{}2,....$ (110)
These integrals were first introduced by Stevenson [18]. Eqn.(107) can be simplified further by using the“gap-equation” as given by eqn.(101-102). One then obtains :
$<~{}{\cal H}_{0}~{}>=I_{1}(M)~{}-~{}3gI_{0}^{2}(M)~{}+~{}\frac{1}{2}~{}m^{2}\sigma^{2}~{}+~{}g~{}\sigma^{4}~{}~{}~{}$ (111)
We thus derive the LO-effective potential of NGAS as given by
$U_{0}(\sigma)=\frac{1}{2}~{}m^{2}\sigma^{2}~{}+~{}g~{}\sigma^{4}~{}+~{}I_{1}(M)~{}-~{}3gI_{0}^{2}(M),~{}~{}~{}$ (112)
where, it is implicitly understood that the “gap-equation”, eqn.(101-102) is to be first solved to obtain ${}M^{2}~{}$as a function of ‘$\sigma$.
IX.4 Equivalence to the Gaussian-Effective Potential
In view of the results derived in the previous subsection , the following remarks/ observations are in order:
(a)  The results contained in eqs. (101-102 ) for the gap equation and the integrals defined by eqs. (108) were first derived by Stevenson [18] in the context of the“Gaussian effective potential (GEP)” for the symmetric ${}g\phi^{4}~{}$ theory and obtained by variational calculation using a Gaussian-trial wave-function. The reproduction of the results of the GEP in ref.[18] in the LO of NGAS demonstrates that the GEP is contained in the NGAS as the leading order approximation.
(b)  In view of the established equivalence with the GEP, all the results obtained in the former approximation, are reproduced in the LO of the NGAS. In particular, the demonstration of non-triviality of the symmetric ${}g\phi^{4}~{}$ theory in the GEP,  being entirely based upon the consequences of renormalization is also reproduced in the LO of NGAS. We elaborate in the following sub-section, some of the properties of the theory concerning the stability and non-triviality of the theory.
(c) It must be emphasized, however, that the current scheme, NGAS is based upon entirely  different  starting  assumptions   and  is much more  general than the GEP, which is obtained solely due to the choice of the AH as given in eqn.(4.3) and that too, in the leading order.
IX.5 Renormalization and realization of the ’non-triviality’ of the theory in the leading order of NGAS
One can carry out the renormalization program (in the LO) by noting that [18] :
(i) the vacuum-state-configuration corresponds to the absolute (global) minimum of ${}U_{0}(\sigma)~{}$, i.e. by solving :
$\frac{dU_{0}}{d\sigma}|_{\sigma_{0}}~{}=~{}0~{};~{}\frac{d^{2}U_{0}}{d\sigma^{2}}|_{\sigma_{0}}~{}>~{}0$ (113)
[…]
$m_{R}^{2}\equiv\frac{d^{2}U_{0}}{d\sigma^{2}}|_{\sigma~{}=~{}\sigma_{0}}~{},~{}~{}~{}~{}$ (114)
(iii)  the LO-renormalized coupling-strength is likewise defined to be:
$g_{R}~{}\equiv~{}\frac{1}{4!}\frac{d^{4}U_{0}}{d\sigma^{4}}|_{\sigma~{}=~{}\sigma_{0}}~{},$ (115)
where ${}\sigma_{0}~{}$ corresponds to the vacuum-configuration as defined by eqn.(111). It is directly verified by minimization of ${}U_{0}(\sigma)~{}$ ( see, eqn.(111)), that the global minimum of the former occurs at ${}\sigma_{0}~{}=~{}0~{}$, which is consistent with $\sigma=0.$ as it should be since the symmetric case is being considered. Next, evaluating eqn.(112) at ${}\sigma_{0}~{}=~{}0~{}$ one gets the renormalized mass :
$m_{R}^{2}~{}=~{}m^{2}~{}+~{}12gI_{0}(m_{R})~{}\equiv~{}M^{2}(~{}g,~{}\sigma^{2}~{}=~{}0),$ (116)
Similarly,  after a straight forward calculation, one obtains ref.[18] the renormalized coupling as given by:
$g_{R}~{}=~{}g[\frac{1-12gI_{-1}(m_{R})}{1+6gI_{-1}(m_{R})}],$ (117)
where again ${}I_{-1}(x)~{}$ is defined as per the general definition given in eqn.(108).
An important consequence of eqn.(114) is that $g$ must be chosen negative: $g<0$, in order that the theory may be renormalized. This is because $I_{0}(m_{R})$ is divergent and the (unmeasurable) bare mass `$m^{\prime}$ of the theory may be infinite but the physical (renormalized) mass $m_{R}$ has to be finite. This version of the theory has therefore, been designated in ref.[18] as precarious ${}g\phi^{4}$ theory.
[…]
IX.6 Emergence of the “Non-Triviality” of the theory
In order to investigate the stability and non-triviality of the theory it is convenient to start with eqs.(101-102), which involve the divergent integral due to the momentum-integration and which, therefore, need a suitable method of subtraction.
Using the subtraction-procedure  devised  by  Stevenson [18], eqn.(115) can be inverted to express ${}g~{}$ in terms of the observable parameters, ${}g_{R}~{}$ and ${}m_{R}~{}$. This leads to two solutions for ${}g~{}$ of which, the physical one is given by :
$g=[-1/6I_{-1}(m_{R})][1+1/2[g_{R}I_{-1}(m_{R})+....]$ (118)
( The other solution is ${}g~{}=~{}-(1/2)g_{R}~{}+~{}O(1/I_{-1}(m_{R})~{}$. This solution can be shown to lead to instability, since the minimum of the EP corresponding to this solution can be shown to lie (infinitely) higher than the minimum corresponding to eqn.(116)).
It may be noted at this point that eqn.(116) implies a viable, stable ${}g\phi^{4}~{}$ theory  when  the  (unobservable)  bare-coupling  becomes negative  but  infinitesimal.
Substituting for ${}g~{}$ as given by eqn.(116), one can solve for the bare-mass, by inverting eqn.(114), after carrying out the subtraction as per the Stevenson-prescription. This leads to the following expression for the bare-mass :
$m^{2}={m_{R}}^{2}+2I_{0}(m_{R})/I_{-1}(m_{R})$
[…]
$U_{0}(\sigma)~{}=~{}U_{min}+\frac{1}{4}t~{}{m_{R}}^{2}\sigma^{2}-({m_{R}}^{4}/128\pi^{2})(t-1)^{2}$
$-({m_{R}}^{4}/64\pi^{2})(t-1)\eta,~{}~{}~{}~{}~{}~{}~{}$ (120)
where
$t~{}=~{}M^{2}(\sigma)/{m_{R}}^{2}~{};~{}~{}\eta\equiv-4\pi^{2}/g_{R}~{},$ (121)
and
$U_{min}=I_{1}(m_{R})-3gI_{0}^{2}(m_{R}).$ (122)
Similarly, the renormalized version of the “gap-equation” is given by :
$(1-\eta)(t-1)-(16\pi^{2}/{m_{R}}^{2})\sigma^{2}~{}=~{}t~{}ln~{}t.$ (123)
[…]
It may be noted that the gap-equation, eqn.(121) is a transcendental equation and its solution exists only when,
$\sigma^{2}~{}\leq~{}\sigma_{min}^{2}~{}\equiv~{}({m_{R}}^{2}/16\pi^{2})[e^{-\eta}~{}+~{}\eta-1]$ (124)
[…]
$<{\cal H}_{0}>_{P}~{}=~{}\frac{1}{2}m^{2}\sigma^{2}~{}+~{}g\sigma^{4}~{}+~{}\bar{I}_{1}~{}+~{}$
$6g\sigma^{2}\bar{I}_{0}~{}+~{}3g{\bar{I}_{0}}^{2}$ (125)
[…]
$U^{P}(\sigma)~{}=~{}\frac{1}{2}m^{2}\sigma^{2}~{}+~{}g\sigma^{4}~{}+~{}$
$~{}\bar{I}_{1}~{}+~{}6g\sigma^{2}\bar{I}_{0}~{}+~{}3g{\bar{I}_{0}}^{2}$ (126)
The renormalized-parameters following from the eqn.(124) are
like-wise computed and denoted by
$\bar{m}_{R}^{2}~{}=~{}d^{2}U^{P}/d\sigma^{2}|_{\sigma^{2}~{}=~{}0}~{};$ (127)
$\bar{g}_{R}~{}=~{}(1/4!)d^{4}U^{P}/d\sigma^{4}|_{\sigma^{2}~{}=~{}0}$ (128)
where ${}\sigma^{2}~{}=~{}0~{}$ is again the location of the global-minimum of ${}U^{P}(\sigma)~{}$, as can be readily verified. In this context, it must be emphasized that the integrals : $\bar{I}_{n}$ occurring in eqn.(124) are $independent$ of $\sigma$.
Next, computing the derivatives of ${}U^{P}(\sigma)~{}$ at the minimum, one obtains the following expressions for the renormalized parameters based upon the perturbative vacuum:
$\bar{m}_{R}^{2}~{}=~{}m^{2}~{}+~{}12g\bar{I}_{0}$ (129)
$\bar{g}_{R}~{}=~{}g$ (130)
[…]
$U^{P}(\sigma)=\frac{1}{2}\bar{m}_{R}^{2}\sigma^{2}+\bar{g}_{R}\sigma^{4}+3\bar{g}_{R}{\bar{I}_{0}}^{2}+\bar{I}_{1}$ (131)
To prevent instability of the theory when renormalised about the perturbative vacuum it, therefore, becomes inescapable that
${}\bar{g}_{R}~{}=~{}g~{}=~{}0~{},$ (132)
[…]
A few remarks/observations are in order, in view of the above results:
(i) We believe that the result in eqn.(130) constitutes perhaps, the most direct demonstration of triviality of symmetric ${}g\phi^{4}~{}$ theory in physical dimensions.
( ii ) At the same time, the result, eqn.(130) also demonstrates that the conclusion of triviality of the theory is an artifact of the $na\ddot{i}ve$ perturbation theory built and renormalized around the free-field vacuum . As demonstrated earlier, the theory renormalized about the NGAS-vacuum leads to a perfectly acceptable, stable and non-trivial ${}g\phi^{4}$- theory ( see, eqs.(118-122) and discussions following ).
( iii ) It may be further pointed out that the ground state of the trivial theory is still unstable as compared to that in the LO of NGAS, i.e.
$U_{min}~{}<<~{}U_{min}^{P},$ (133)
which is readily established by referring to eqn.(129) (with ${}\sigma~{}=~{}0~{}$) and eqn.(120).
This completes our results and discussions regarding the stability and the triviality of ${}g\phi^{4}$- theory in the context of NGAS in the LO.
In the next subsection, we discuss the structure of the interacting vacuum .
IX.7 The Underlying Bogoliubov-Valatin Transformation
The actual/physical vacuum state in presence of interaction is approximated in the LO of NGAS by the state: $|~{}vac>$ which is the lowest energy state of ${}H_{0}~{}$. The structure and properties of this state can be inferred from studying the quantum-canonical transformation (i.e. Bogoliubov-Valatin transformation, ref.[46]) connecting the interacting vacuum state (IVS) with the free field vacuum ( FFV ) state.
[…]
$\phi(\vec{x},t)=\sigma+\int\frac{d^{3}\vec{k}}{\Omega_{k}(m)}[a(\vec{k})e^{-ikx}+a^{\dagger}(\vec{k})e^{ikx}],~{}~{}~{}~{}~{}~{}~{}$ (134)
where, now
$\Omega_{k}(m)~{}\equiv~{}2~{}(2\pi)^{3}~{}\omega_{k}(m)~{};$ (135)
$~{}{k}^{0}~{}=~{}\omega_{k}(m)~{}\equiv~{}\sqrt{|\vec{k}|^{2}~{}+~{}m^{2}},$ (136)
corresponding to the propagation of the free-field quanta satisfying the mass-shell condition : ${}{k^{0}}^{2}~{}-~{}|\vec{k}|^{2}~{}=~{}m^{2}~{}$. The free-field operators satisfy the standard commutation relations :
$[a(\vec{k}),~{}a^{\dagger}(\vec{q})]~{}=~{}\Omega_{k}(m)~{}\delta^{3}(\vec{k}-\vec{q}).$ (137)
Comparision of eqs.(90) and (135) implies that the modified operators :
$B(\vec{k})~{}\equiv~{}\frac{b(\vec{k})}{\sqrt{\Omega_{k}(M)}}$ (138)
$A(\vec{k})~{}\equiv~{}\frac{a(\vec{k})}{\sqrt{\Omega_{k}(m)}}$ (139)
satisfy identical commutation relations :
$[B(\vec{k}),B^{\dagger}(\vec{q})]~{}=~{}\delta^{3}(\vec{k}-\vec{q})~{}=~{}[A(\vec{k}),~{}A^{\dagger}(\vec{q}))].$ (140)
It follows, therefore, that the two sets of operators must be connected through Boguliobov-transformation ref.[46] given by :
$B(\vec{k})~{}~{}=~{}cosh(\alpha_{k})A(\vec{k})~{}-~{}sinh(\alpha_{k})A^{\dagger}(-\vec{k})$ (141a)
$B^{\dagger}(\vec{k})=~{}cosh(\alpha_{k})A^{\dagger}(\vec{k})~{}-~{}sinh(\alpha_{k})A(-\vec{k}),$ (141b)
whereas the inverse transformation is given by :
$A(\vec{k})~{}~{}=~{}~{}cosh(\alpha_{k})B(\vec{k})~{}+~{}sinh(\alpha_{k})B^{\dagger}(-\vec{k})$ (142a)
$A^{\dagger}(\vec{k})~{}=~{}cosh(\alpha_{k})B^{\dagger}(\vec{k})~{}+~{}sinh(\alpha_{k})B(-\vec{k}).$ (142b)
In the above, ${}\alpha_{k}~{}=~{}f(|\vec{k}|)~{}$ , is apriori an arbitrary real function of ${}|\vec{k}|~{}$, i.e.
$\alpha_{-\vec{k}}~{}=~{}\alpha_{\vec{k}}~{}=~{}\alpha_{\vec{k}}^{*}.$ (143)
However, eqs.( 90 ), (132) and (139) considered together further imply that
$exp~{}(2\alpha_{k})~{}=~{}\frac{\omega_{k}(M)}{\omega_{k}(m)}~{}=~{}\frac{\sqrt{|\underline{k}|^{2}+M^{2}}}{\sqrt{|\underline{k}|^{2}+m^{2}}}$ (144)
To show this, consider eqs.(90) and (132) at ${}t~{}=~{}0~{}$, which can be written as:
$\phi(\vec{x},t)=\sigma+\int\frac{d^{3}\vec{k}}{\sqrt{\Omega_{k}(m)}}[~{}A(\vec{k})+A^{\dagger}(-\vec{k})]e^{i\vec{k}.\vec{x}}$ (145a)
$=\sigma+\int\frac{d^{3}\vec{k}}{\sqrt{\Omega_{k}(M)}}[B(\vec{k})+B^{\dagger}(-\vec{k})]e^{i\vec{k}.\vec{x}},$ (145b)
which implies that :
$\{B(\vec{k})+B^{\dagger}(-\vec{k})\}=\sqrt{\frac{\Omega_{k}(M)}{\Omega_{k}(m)}}\{A(\vec{k})+A^{\dagger}(-\vec{k})\}$ (146)
However, from eqn.(139) it follows that
$\{B(\vec{k})~{}+~{}B^{\dagger}(-\vec{k})\}~{}=~{}exp~{}(\alpha_{k})\{~{}A(\vec{k})~{}+~{}A^{\dagger}(-\vec{k})~{}\}$ (147)
thus leading to the desired result, eqn.(142).
The significant physical results that follow from the above equations are discussed in the following subsection.
IX.8 Structure and Stability of the Effective Vacuum
To obtain the information regarding the particle-content and other features of the IVS it is instructive to first compute the number-density of the free-field-quanta residing in the IVS. To this end let us note that the free-field-number operator is given by the standard expression:
$N~{}\equiv~{}\int{\frac{d^{3}\vec{k}}{\Omega_{k}(m)}a^{\dagger}(\vec{k})a(\vec{k})}$
$~{}=~{}\int{d^{3}\vec{k}}~{}A^{\dagger}(\vec{k})~{}A(\vec{k})$ (148)
Hence the desired number density of the free-field quanta in the IVS defined as:
$n(\vec{k})~{}\equiv~{}<~{}vac~{}|~{}\frac{A^{\dagger}(\vec{k})A(\vec{k})}{v}~{}|~{}vac>~{},$ (149)
where ${}v~{}\equiv~{}$ spatial-volume of quantization ${}\equiv~{}\int d^{3}\vec{x}$ can be easily evaluated, using eqs.(140) and given by
$n(\vec{k})~{}=~{}\frac{sinh^{2}(\alpha^{vac}(\vec{k}))}{(2\pi)^{3}}$ (150)
where ${}\alpha^{vac}(\vec{k})~{}$ is given by eqn.(142) evaluated for ${}M~{}\equiv~{}m_{R}~{}$=  free-particle-mass renormalized  about the IVS, ${}|vac>$  (It may be recalled that ${}M(\sigma~{}=~{}0)~{}=~{}m_{R}~{}$ and ${}\sigma~{}=~{}0~{}$ define the IVS ). This leads finally to the expression :
$n(\vec{k})~{}=~{}(\frac{1}{32\pi^{3}})~{}[~{}\frac{\omega_{k}(m)}{\omega_{k}(m_{R})}~{}+~{}\frac{\omega_{k}(m_{R})}{\omega_{k}(m)}~{}-~{}2~{}]~{}.$ (151)
[…]
$\lim_{\Lambda\rightarrow\infty}~{}(\frac{n(\vec{k})}{n(\vec{0})})~{}\equiv~{}\rho(\vec{k})~{}=~{}(1~{}+~{}\frac{|\vec{k}|^{2}}{m_{R}^{2}})^{-\frac{1}{2}}~{},$ (152)
where ${}n(\vec{0})~{}=~{}n(\vec{k})|_{max}~{}=~{}(\frac{1}{32\pi^{3}})(\frac{m}{m_{R}})~{}$, is the maximum value of ${}n(\vec{k})~{}$, occuring at ${}\vec{k}~{}=~{}0$.
Equation (150)provides direct physical content for the non-trivial structure of the IVS representing a condensate of off-shell correlated particle-pairs. The situation is analogous to the structure of the physical vacuum state in case of the hard-sphere Bose-gas [43] and superfluidity . It is therefore , plausible that eqn.(150) might lead to interesting consequences for ${}T~{}\neq~{}0~{}$, as happens in the case of the super-fluid and the hard-sphere Bose-gas.
IX.9 Section-IX : Summary
We have demonstrated that the physical properties of the $\phi^{4}$ field theory depend critically on the structure of the physical vacuum state. The interacting vacuum state leads to a non-trivial and stable theory with finite renormalized parameters whereas, the physics developed around the non-interacting (‘perturbative’) vacuum becomes either unstable or trivial.
X Perturbation theory for arbitrary coupling Strength in NGAS - motivation and scope
We have discussed the key features of the standard formulation of perturbation theory (SFPT) in Section-I while surveying different approximation methods and noted its merits and the limitations. In this Section, we propose an improved perturbation theory, herein after designated as “Mean Field PerturbationTheory”(MFPT) which is based on NGAS and aimed at overcoming the limitations of SFPT. (The nomenclature will be explained in the sections to follow.) Some clues of how to achieve this objective have been provided at the end of Section-VIII and these will be further developed in this Section. In particular, we will explore how to preserve in MFPT, the nonlinearity and the analytic properties of the system-Hamiltonian and other “non-perturbative” effects , which are not achievable in SFPT.
However, there are certain features, tools and techniques used in SFPT which may be needed also in MFPT. In the following Section we consider those particular ’tools’ and ’techniques’ which may be relevant for this purpose.
XI General features of the ‘standard’ formulation of perturbation theory(SFPT)
[…]
$E(g)=\mathcal{E}_{0}+\sum\limits_{k=1}^{\infty}{g^{k}\mathcal{E}_{k}\equiv\mathcal{E}_{0}(g)+\Delta{\mathcal{E}(g)}},$
where, the first term $\mathcal{E}_{0}$ represents the contribution known from the non-interacting theory while $\mathcal{E}_{k}$ denotes the perturbative correction (PC) at the $k$-th order. In order that the PC here make sense it is ab initio necessary that the coupling strength, $g$ be sufficiently small, i.e. ${}g\ll 1$.
[…]
$|\mathcal{E}_{k}|\sim c^{k}\Gamma(ak+b),for~{}k\gg~{}1,$ (153)
where $a,b,c$ are constants depending upon the particular theory under consideration. Moreover, the asymptotic-nature of the PS manifests through the following property [107,108]:
$\lim_{g\rightarrow 0_{+}}(~{}E(g)-\sum\limits_{k=0}^{N}g^{k}\mathcal{E}_{k}~{})/~{}g^{N+1}=\mathcal{E}_{N+1}.$ (154)
In that context, historically the systematic investigations of large-order behavior of (renormalized) perturbation expansion were primarily established by studies [88] of systems with anharmonic interactions (AHI) in quantum mechanics (QM) and through subsequent investigations [107] in other systems.
It is now established, (see, e.g.[98,107,108]) that generalized-factorial growth of perturbation correction is a general feature of perturbation theory for almost all cases of physical interest and further that this behavior is primarily due to the factorial growth of Feynman diagrams/ intermediate states contributing at large orders irrespective of the theory under consideration.
Therefore, It becomes foremost necessary, to assign a meaning to the asymptotic-perturbation theory and investigate whether the latter can be “summed” in some sense to yield a finite result. We will focus on two main methods to achieve this goal, namely the method of “Optimal Truncation of the Asymptotic Series” and the “method of Borel-summation”.These are discussed in some detail in the next subsection.
XI.1 Making sense of divergent, asymptotic series: Optimal truncation
This method utilizes the property [107,108], that the initial terms of an AS continue to decrease in magnitude till the term of the least magnitude(TLM) is reached, beyond which the subsequent terms exhibit monotonic increase. Hence, the AS can be reasonably approximated by truncation at the TLM.
XI.2 Making Sense of Divergent, Asymptotic Series: Borel-Resummation
The method of Borel-summation is considered [107,108] the standard procedure for constructing the analytic function, $E(g)$ from its divergent, asymptotic series for $g\rightarrow{0_{+}}$. Indeed, the ‘sum’ of the divergent, asymptotic PS is now customarily defined [107,108] by its Borel-sum when the latter exists. However, Borel-summability holds under certain restrictive conditions imposed on the original PS . In particular, it is a necessary condition [107,108] that the successive terms in eqn.(1) must alternate in sign at large-order, $k\gg 1$ to ensure Borel-summability.
This requirement exposes further limitations of SFPT , since it is found that several physically important theories, such as the case of the quartic double-well oscillator (QDWO) and other cases exhibiting degenerate vacua /ground states, fail [95] this necessary condition for Borel-summability.
Furthermore, in the context of SFPT, both the above methods are restricted to the perturbative domain characterized by small values ( compared to unity) of the coupling $g$.
In view of the inadequacies of SFPT mentioned in Section-II and the previous section, it is perhaps imperative to explore whether perturbation theory could at all be liberated from the limitations of the small-coupling-power-series-expansion. The following Sections devoted to such an exploration in some detail.
XII General Features of the ‘Mean Field Perturbation Theory’ (MFPT) in NGAS
One of the main motivations for proposing the NGAS as described earlier, is the possibility of construction of an improved perturbation theory which could be convergent/summable for all allowed values of the coupling strength, $g$.
Working in the standard Hamiltonian formalism, we can construct the improved perturbation theory in which the system-Hamiltonian, $H(g)$ can be split in to an exactly solvable dominant part $H_{0}$ and a sub-dominant perturbation $H^{\prime}$, i.e.
$H(g)=H_{0}+H^{\prime}.$ (155)
Then, for the realization of a perturbative-framework for arbitrary values of the coupling $g$, it may be sufficient to fulfill the following two conditions:
(a) both $H_{0}$ and $H^{\prime}$ must depend non-trivially upon the coupling strength $g$, i.e. $H_{0}=H_{0}(g)$ and $H^{\prime}=H^{\prime}(g)$ such that the dominant $g$-dependence resides in $H_{0}(g)$, and further that (b) the contribution of the perturbing Hamiltonian $H^{\prime}(g)$ remains sub-dominant for arbitrary value of $g$.
As discussed earlier ( Section-VI), a practical implementation of the above conditions can be achieved [82,104] through the following steps:
(i) choose $H_{0}$ to depend upon a suitable set of free parameters ${{\alpha_{i}}}$: $H_{0}=H_{0}(\{\alpha_{i}\})$,
(ii) impose the constraint that
$\langle\phi_{n}|H(g)|\phi_{n}\rangle~{}=~{}\langle\phi_{n}|H_{0}(\{\alpha_{i}\})|\phi_{n}\rangle$ (156)
where,$|\phi_{n}\rangle$ is defined by the eigen-value equation for $H_{0}$ given by
$H_{0}|\phi_{n}\rangle~{}=~{}E_{0}^{n}|\phi_{n}\rangle,$ (157)
with ``$n"$ denoting the spectral-label,
(iii)finally, determine the parameters $\{\alpha_{i}\}$ by variational minimization of $\langle H(g)\rangle$ and subsequent imposition of the constraint given by eqn.(154).
By this simple procedure, the non-linear $g$-dependence of the system-Hamiltonian, $H(g)$ is fed back into the reference Hamiltonian:
$H_{0}(\{\alpha_{i}\})\rightarrow~{}H_{0}(g)$ (158)
Solving the eigen-value equation (eqn.155) for $H_{0}(g)$, then leads to [82,104] the results for $E_{0}(g)$ which, constitute the leading-order (LO) approximation. These LO-results are found [82,104] to be quite accurate uniformly over all spectral-labels ‘$n$’and for arbitrary physical values of $g$. It may be noted that this procedure is non-perturbative, self-consistent and preserves the analytic-structure of $E(g)$ in the $g$-plane [82,104].
The above procedure has already been discussed under Section-VI but succinctly recapitulated here for subsequent discussions.
Thus, having achieved the incorporation of the major $g$-dependence at the LO, the task remains to build a perturbation theory about the approximating-Hamiltonian (AH), $H_{0}$ and systematically improve the accuracy still further by computing order-by-order, the residual corrections as described in the next subsection.
XII.1 Definition and Implementation
A new formulation of perturbation expansion [110,111], which we designate as “Mean-Field-Perturbation Theory (MFPT)”, is naturally suggested in view of the above considerations by defining the perturbation about the AH as:
$H^{\prime}(g,n)\equiv H(g)-H_{0}(g,n).$ (159)
An immediate consequence of this prescription , eqn.(157) is the following result:
$\langle\phi_{n}|H^{\prime}(g,n)|\phi_{n}\rangle=0,$ (160)
for all ``$n"$ and ``$g"$, by virtue of eqns.(154) and (156). This eqn. (158) ensures that the perturbation correction remains sub-dominant, in the sense of quantum average , for arbitrary values of `$g^{\prime}$ and `$n^{\prime}$ as demanded (see condition-(b) mentioned above, i.e.
$\langle H^{\prime}(g)\rangle\ll\langle H_{0}(g)\rangle\equiv{E}^{n}_{0}(g).$ (161)
(Here,the quantum average of an operator A is defined as: $\langle A\rangle\equiv\langle\phi_{n}|A|\phi_{n}\rangle)$.
Furthermore, eqn. (158) has a direct consequence that the first-order perturbation correction in the Rayleigh-Schrö dinger-perturbation series (RSPS) vanishes identically for all `$n^{\prime}$ and `$g^{\prime}$:
${E}^{n}_{1}\equiv\langle\phi_{n}|H^{\prime}|\phi_{n}\rangle=0,$ (162)
It is shown later that these two general outcomes of eqn.(158) have important bearing on the properties of the perturbation-series (PS) in MFPT, which is defined analogous to eqn.(1) as follows:
$E(g)=E_{0}(g)+\sum_{k=1}^{\infty}E_{k}(g)\equiv E_{0}(g)+\Delta E(g).$ (163)
(hence-forth, we do not display the $n$-dependence of the various quantities, for notational convenience)
For the study of convergence properties of the above series, one needs to compute the energy-corrections $E_{k}(g)$ to arbitrary order `$k^{\prime}$. However, since the above series is not a power- series-expansion in $g$, we resort to the well-known [109] recipe of introducing an auxiliary, dummy parameter denoted as $\eta$, for generating a power series in this parameter (chosen real) and project out the `$k^{\prime}$-th order correction, $E_{k}(g)$ by the following procedure: consider an associated-Hamiltonian (ASH), $\bar{H}$ given by
$\bar{H}\equiv H_{0}+\eta H^{\prime},$ (164)
and the corresponding eigenvalue equation:
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}\bar{H}(\eta,g)|\bar{\psi}(\eta,g)\rangle=\bar{E}(\eta,g)|\bar{\psi}(\eta,g)\rangle,$ (165)
such that $\bar{E}(\eta,g)$ can be expanded as a formal power series in $\eta$ as follows:
${}~{}~{}~{}~{}~{}~{}~{}\bar{E}(\eta,g)=\sum_{k=0}^{\infty}\eta^{k}E_{k}(g)$ (166)
The `$k^{\prime}$-th perturbation correction $E_{k}(g)$ being the same as appearing in eqn.(161), can then be identified from the above equation as the coefficient of $\eta^{k}$ before setting the limit, $\eta\rightarrow 1$.
It may be emphasized that the above procedure is merely an intermediate book-keeping device to project out the $E_{k}(g)$ appearing in eqn. (161). Apart from achieving this goal, the above formalism plays no other role here. In particular, the final results are independent of the dummy-variable $\eta$ since, by construction, $H(g)=\bar{H}(\eta=1,g)$, $E(g)=\bar{E}(\eta=1,g)$, etc.
We determine the corrections $E_{k}(g)$ to arbitrary order `$k^{\prime}$ in MFPT, using the recursion relations derived from the application of the “hyper-virial theorem (HVT)” and the “Feynman-Hellman theorem (FHT)” to the Hamiltonian given by eqn.(163). This method is distinguished from other methods in generating exact values for perturbation corrections to an arbitrary order whereas other methods are afflicted by inevitable errors such as neglecting sub-asymptotic corrections, truncation-errors etc.(Moreover, the ease of implementation of recursive evaluation through widely available symbolic-computation, is an additional practical advantage of this method.) Here, we follow procedure analogous to that in ref.[9] used in the context of SFPT for the QAHO, with due generalization for extending the application to MFPT as outlined below in the next section.
XIII Application of MFPT to Anharmonic -interactions(AHI) in the Harmonic Approximation
XIII.1 Tools: The Feynman-Hellman Theorem (FHT) and the Hyper-Virial Theorem (HVT)
In many cases of physical interest, the simultaneous application of the Feynman-Hellman Theorem and the hyper-virial theorem, renders the computation of perturbation corrections $E_{k}$ much easier without the necessity of summing infinite (sub)-series at each order. Furthermore, each term: $E_{k}$ can be precisely computed (with zero error) as inevitably occurs due to finite truncation of the conventional perturbation series at each order. Moreover, precise computation of $E_{k}$ to an arbitrary high-value of ‘$k$’ can be exactly computed through (system-specific-) recursion relations. For the AHI, the program was carried out by Swenson and Danforth [81] and elaborated in texts e.g. by Fernandez[9]. We discuss these theorems in turn, as follows.
XIII.1.1 The Hyper-virial Theorem(HVT)
The content of the Hyper-Virial-Theorem (HVT) is given by the following equation , see, e.g.ref.[9]:
$\langle\psi|\left[\hat{H},\hat{W}\right]|\psi\rangle~{}=~{}0,$ (167)
where,$\hat{H}$ is the (hermitian) Hamiltonian of the system and $[\hat{A},\hat{B}]$ stands for the commutator of operators, $\hat{A}$ and $\hat{B}$ and $\hat{W}\equiv\hat{W}{(\hat{x},\hat{p})}$ is any arbitrary, time independent operator function of $\hat{x}$ and $\hat{p}$ in the relevant Hilbert space. Here, we work in the co-ordinate space- representation such that $\hat{p}~{}=~{}-i\frac{\partial}{\partial{x}}$ and $\hat{x}~{}=~{}x$ are the momentum and position operators.
The above eqn.(165) is a consequence of the Energy-eigenvalue equation(and its hermitian conjugate) as appropriate for the time-independent ( ’Stationary’) situation and can be verified directly using :
$\hat{H}|\psi\rangle~{}=~{}E|\psi\rangle,\langle\psi|\hat{H}~{}=~{}E\langle\psi|$ (168)
We consider next the Hellmann-Feynman Theorem.
XIII.1.2 The Hellmann-Feynman Theorem(HFT)
Consider when the Hamiltonian, ${H}$ of the system depends on a parameter $\lambda$ , apart from the dynamic-variables of the system under consideration. (This situation is usually the case since $\lambda$ could be the system- charge, mass, a coupling-strength of interaction or even a dummy parameter introduced to define perturbation theory etc.) Then it follows that the eigenvalues and eigenfunctions of ${H}$ will also depend on that parameter.
The Hellmann-Feynman-Theorem (HFT) applies to this situation and leads to profound results in perturbation theories without the use of wave-functions, see e.g. ref.[9] .The theorem states that:
The partial derivative of the total energy with respect to an parameter occurring in the Hamiltonian of the system is equal to the expectation value of the partial derivative of the Hamiltonian with respect to the same parameter, as given by the following equation:
$\langle\psi_{n}|\left(\frac{\partial H}{\partial\lambda}\right)|\psi_{n}\rangle~{}=~{}\frac{\partial E_{n}}{\partial\lambda}.$ (169)
The proof (see,e.g.[9]) of the theorem is a straight forward application of the eigenvalue equation and the hermitian-property of the Hamiltonian for a stationary system , as given by:
$(H-E_{n})|\psi_{n}\rangle~{}=~{}0~{}=~{}\langle\psi_{n}|(H-E_{n}),$ (170)
which yields eqn.(167) on partial differentiation with respect to $\lambda$ if $|\psi_{n}\rangle$ is normalized to unity.
( In the above context, it may be relevant to remark that: (i) there are generalizations of the (FHT) to its ‘off-diagonal’ version, see ref.[106] and that (ii) the off-diagonal-FHT finds important applications e.g. in reproducing the RS-Perturbation theory without the use of the unperturbed wave-function).
XIII.2 Tools : The Method of Swenson and Danforth
The method of Swenson and Danforth [ 81], utilizes the hyper-virial and the Hellmann-Feynman theorems to derive the perturbation-correction for the anharmonic oscillators in the SFT-framework. A lucid description of the same is provided in ref.[9] , which we closely follow in this subsection and present an overview of the method.
To start with, we consider the Hamiltonian for a non-relativistic, stationary system in one space-dimension given by:   $\hat{H}=\hat{p}^{2}{/}{2m}+V(\hat{x})$, and use the co-ordinate representation such that $\hat{x}=x$ and $\hat{p}=-i\hbar\partial{/}\partial x$.
In the next step, we explore the consequence of the HVT by choosing a suitable linear operator, $\hat{W}$ depending up on $\hat{x}$ and $\hat{p}$ Following ref. [9] , the following choice is made :
$\hat{W}~{}=~{}f(x)\hat{p}+g(x),$ (171)
where, $f(x)$ and $g(x)$ are arbitrary, ’regular’ (differentiable) functions to start with but allowing to be chosen conveniently for the system under study. Considering the one-dimensional Hamiltonian operator as given above, i.e.   $\hat{H}=\hat{p}^{2}{/}{2m}+V(\hat{x})$,   and using eqn.(169) in eqn.(165), one obtains:
$\langle\left[H,fp+g\right]\rangle~{}=~{}\langle\left[H,fp\right]\rangle+\langle\left[H,g\right]\rangle~{}=~{}0,$ (172)
where, $\langle~{}\hat{A}~{}\rangle$ denotes   the ${}expectation~{}value~{}of~{}the~{}operator~{}\hat{A}~{}in~{}the~{}state,~{}\rvert\psi\rangle~{}$. This equation can be expressed as the sum of two terms i.e. ${}A+B~{}\equiv~{}0$, where ${}A=\langle\left[H,fp\right]\rangle~{}$ and ${}B=\langle\left[H,g\right]\rangle~{}$.
Using the basic commutation relation,   $[\hat{x},\hat{p}]=-i\hbar$   and the result: $[\hat{p},f(x)]=i\hbar f^{\prime}$ (where denotes derivative with respect to ${}x~{}$) and employing the standard manipulation techniques for $commutators$, one obtains the following result for $A$:
$A~{}=~{}\left(\frac{-\hbar^{2}}{2m}\right)\langle{f^{\prime\prime}}p\rangle-\left(\frac{i{\hbar}}{m}\right)\langle{f^{\prime}}p^{2}\rangle+(i{\hbar})\langle f{V^{\prime}}\rangle.$ (173)
Similarly, the $B$ term can be evaluated as
$B~{}=~{}\langle\left(\frac{-\hbar^{2}}{2m}\right){g^{\prime\prime}}\rangle-\left(\frac{i{\hbar}}{m}\right)\langle{g^{\prime}}p\rangle.$ (174)
The substitution of $A$ and $B$ in the defining equation,
$0~{}=~{}A+B\equiv\langle\left[H,W\right]\rangle$ (175)
then yields the result
$\left(\frac{-\hbar^{2}}{2m}\right)\langle{f^{\prime^{\prime}}}p\rangle-\left(\frac{i{\hbar}}{m}\right)\langle{f^{\prime}}p^{2}\rangle+(i{\hbar})\langle f{V^{\prime}}\rangle+\left(\frac{-\hbar^{2}}{2m}\right)\langle{g^{\prime\prime}}\rangle-\left(\frac{i{\hbar}}{m}\right)\langle{g^{\prime}}p\rangle~{}=~{}0.$ (176)
Next, since $f$ and $g$ are arbitrary regular functions, one can choose $g$ in terms of $f$ such that term linear in $\hat{p}$ vanishes. This is achieved by imposing the following constraint:
$\left(\frac{-\hbar^{2}}{2m}\right)\langle{f^{\prime\prime}}p\rangle+\left(\frac{-i{\hbar}}{m}\right)\langle{g^{\prime}}p\rangle~{}=~{}0.$ (177)
The constraint-eqn.(175) is thus satisfied by the following choice of the g-function:
$g~{}=~{}\left(\frac{1}{2}\right)\left[p,f\right]~{}=~{}\left(\frac{i{\hbar}}{2}\right){f^{\prime}}.$ (178)
Substitution of eqn.(176) in eqn.(174) then leads to
$\left(\frac{-i{\hbar}}{m}\right)\langle{f^{\prime}}p^{2}\rangle+(i{\hbar})\langle f{V^{\prime}}\rangle-(i{\hbar})\left(\frac{\hbar^{2}}{4m}\right)\langle{f^{\prime\prime\prime}}\rangle~{}=~{}0.$ (179)
The insertion of
$\hat{p}^{2}|\psi\rangle~{}=~{}\{2m(E-V)\}\psi\rangle,$ (180)
in eqn.(177) and simplifying, then yields:
$02E\langle{f^{\prime}}\rangle-2\langle{f^{\prime}}V\rangle-\langle f{V^{\prime}}\rangle+\left(\frac{\hbar^{2}}{4m}\right)\langle{f^{\prime\prime\prime}}\rangle~{}=~{}$ 0 . (181)
In the above equation, the notation is as follows: $E$ = energy eigenvalue; $f=f(x)$ is an arbitrary differentiable function which can be conveniently chosen for the particular problem at hand,“prime(s)” denote differentiation and $\langle A\rangle\equiv\langle\psi|A|\psi\rangle$ for an operator A, with $|\psi\rangle$ denoting the (normalized) eigenfunction of $H$.
It may be noted here that: (a) alternative derivation of equation (179) can be achieved, e.g. [9] by using the Jacobi identity for commutators. (b) This equation along with use of the Hellmann-Feynman Theorem, will be shown (later in this Section) to lead to recursive derivation of exact perturbation-corrections at ‘arbitrary’ order. (c) The yet-arbitrary function ‘$f$’ can be chosen according to the specific functional form for the potential-function: $V(x)$.
In the following subsection, we illustrate the SD-Method by applying it to the case of the general an-harmonic interaction in one-space dimension.
XIII.3 The Swenson-Danforth Method (SDM) for the An-Harmonic Interactions (AHI)
Consider the Hamiltonian(dimensionless units used):
$H~{}=~{}\frac{1}{2}p^{2}+V$ (182)
where
$V~{}=~{}\frac{1}{2}x^{2}+\lambda x^{2K},K=1,2,3,....$ (183)
Rewriting eqn.(179) in these notations leads to:
$02E\langle{f^{\prime}}\rangle-2\langle{f^{\prime}}V\rangle-\langle f{V^{\prime}}\rangle+(\frac{1}{4})\langle{f^{\prime\prime\prime}}\rangle~{}=~{}$ 0 . (184)
For the even an-harmonic oscillator as defined in eqn.(181), it is suggested that we choose:
$f(x)~{}=~{}x^{2j+1}$ (185)
where’$j$’= 0,1,2,3,4….. Computing ${f^{\prime}}$, ${V^{\prime}}$, ${f^{\prime\prime\prime}}$, etc, substituting into eqn.(182) and simplifying , one obtains:
$2E(2j+1)\langle x^{2j}\rangle-2(2j+1)\langle x^{2j}(\frac{1}{2}x^{2}+\lambda x^{2K})\rangle$
$-\langle x^{2j+1}(x+\lambda(2k)x^{2K-1})\rangle$
$+(\frac{1}{4})(2j+1)(2j)(2j-1)\langle x^{2j-2}\rangle$ $=$ $0.~{}~{}~{}~{}~{}~{}~{}~{}$ (186)
Denoting,
$\langle x^{2j}\rangle\equiv X(j),$ (187)
and substituting eqn.(185) into eqn. (184), yields
$2E(2j+1)X(j)-2(2j+1)(\frac{1}{2}X(j+1)$
$+\lambda X(j+K))-\{X(j+1)+\lambda(2k)X(j+K)\}$
$+(\frac{1}{4})(2j)(4j^{2}-1)X(j-1)$ $=$ $0.~{}~{}~{}~{}~{}~{}~{}$ (188)
Rearranging eqn.(185) one finally gets the desired recurrence relation for the expectation values of $X(j)$
$X(j+1)$ $=$ $\frac{1}{2(j+1)}\{2(2j+1)EX(j)$ (189)
$+\frac{j(4j^{2}-1)}{2}X(j-1)$
$-2\lambda(2j+K+1)X(j+K)\}$
As a check of eqn.(187) , consider the special case when $\lambda=0$. Then one recovers the $Hamiltonian$ of the simple-harmonic-oscillator (SHO), given by
$H\rightarrow H_{SHO}\equiv\frac{1}{2}p^{2}+V_{0}~{};~{}~{}V_{0}$ $=$ $\frac{1}{2}x^{2}~{}~{}~{}~{}~{}$ (190)
Since this case is exactly solvable, we know the eigenvalues ,
$E\rightarrow{\cal E}\equiv E_{SHO}~{}=~{}(n+\frac{1}{2})\equiv\xi$ (191)
where , $n=0,1,2,3,......$. (Henceforth, we suppress the label ${}^{\prime}n^{\prime}$ in $E$ and ${\cal E}$ for notational convenience.) Denote the eigenstates of $H_{SHO}$ by
$|\phi_{n}\rangle\equiv~{}{\lim}_{\lambda\rightarrow 0}~{}~{}|\psi_{n}\rangle.$ (192)
Also, let us denote by $Y(j)$:
$Y(j)\equiv\langle\phi_{n}|x^{2j}|\phi_{n}\rangle$ (193)
Then from eqn.(187), substituting $\lambda=0$ and using eqns.(190-191) we get
$Y(j+1)=\left\{(\frac{2j+1}{j+1})\right\}{\cal E}Y(j)+\left\{\frac{j(4j^{2}-1)}{4(j+1)}\right\}Y(j-1)~{}~{}~{}~{}$ (194)
Recursive solution for $Y(j)$; $j$=0,1,2,…. can be obtained from eqn.(192)in terms of the known eigenvalues ${\cal E}$ Noting that: $Y(0)~{}=~{}1$ and $Y(j)\equiv 0$ for $j<0$ (by choice), we get
$Y(1)={\cal E}Y(0)={\cal E}\equiv\langle x^{2}\rangle_{SHO},$
$Y(2)=(\frac{3}{2}){\cal E}Y(1)+(\frac{3}{8})Y(0),$
$=(\frac{3}{2}){\cal E}^{2}+(\frac{3}{8})\equiv\langle x^{4}\rangle_{SHO},$
$Y(3)=(\frac{5}{3}){\cal E}Y(2)+\frac{5}{2}Y(1)\equiv\langle x^{6}\rangle_{SHO}~{}~{},$ (195)
etc. , which are the standard results (see, eqns.(27a-27d)), obtained by other methods, e.g. the use of ’ladder- operators’ for the SHO.
Let us next consider the case when $\lambda\neq 0$ in eqs.(180-181). Then, the Hamiltonian, (eqn.(180)) corresponds to the case of the anharmonic oscillator with anharmonicity = $2K$ (=4,6,8,….). Now of course, $E$ is not known analytically. However, if the standard formulation of perturbation theory (SFPT) is applied (with $\lambda$ considered as the expansion parameter) then one proceeds by expanding the observables appearing in the ‘Master Formula’ (eqn.(187) as follows:
$E\equiv E(\lambda)=\sum_{p=0}^{\infty}E_{p}\lambda^{p},~{}~{}~{}X(j,\lambda)\equiv\sum_{i=0}^{\infty}\lambda^{i}X(j,i).$ (196)
with $E_{0}\equiv{\cal E}=\xi\equiv(n+\frac{1}{2})$ = SHO-energy levels.
[…]
$\langle\frac{\partial H}{\partial\lambda}\rangle~{}=~{}\langle x^{2K}\rangle\equiv X(K,\lambda)$
$~{}=~{}\sum_{p=0}^{\infty}\lambda^{p}X(K,p).$ (197)
and
$\frac{\partial E}{\partial\lambda}~{}=~{}\sum_{p=1}^{\infty}\lambda^{p-1}(pE_{p})$ (198)
Equating the two expressions, eqns.(195-196) as required by HFT ( eqn.(167) ) then leads to :
$\sum_{p=1}^{\infty}\lambda^{p-1}\{X(K,p-1)-pE_{p}\}~{}=~{}0.$ (199)
Hence, equating the co-efficient of each power of $\lambda$ to zero in eqn.(197) yields:
$E_{p}~{}=~{}(\frac{1}{p})X(K,p-1)$ (200)
This eqn.(198) thus relates the perturbation correction at order-$p$ in terms of the $X(K,p-1)$ , which can be evaluated recursively by the recurrence relation given by eqn.(187),again by expanding the relevant observables in power-series in $\lambda$ as given by eqn.(194) and reproduced below:
[…]
where $X(j,i)$ $\sim$ Taylor-coefficients in power-series expansion of $X(j,\lambda)$, i.e.
$X(j,i)~{}=~{}(\frac{1}{i!})\frac{\partial^{i}}{\partial\lambda^{i}}X(j,\lambda)|_{\lambda=0};$ (201)
Next, consider the power-series expansion for the product: $X(j,\lambda)E(\lambda)$ given by
$X(j,\lambda)E(\lambda)~{}=~{}\sum_{i=0}^{\infty}f_{i}\lambda^{i}$ (202)
Then, from the rules of power-series multiplication , the coefficients $f_{i}$ get determined in terms of $E_{m}$ and $X(j,i)$ given by
$f_{i}~{}=~{}\sum_{m=0}^{i}E_{m}X(j,i-m)$ (203)
After substitution of the above equations in eqn.(187), leads to
$\sum_{i=0}^{\infty}\lambda^{i}X(j+1,i)$ $=$ $\left(\frac{2j+1}{j+1}\right)\sum_{i=0}^{\infty}\lambda^{i}\{\sum_{m=0}^{i}E_{m}X(j,i-m)\}$
+$+\left\{\frac{j(4j^{2}-1)}{4(j+1)}\right\}\sum_{i=0}^{\infty}\lambda^{i}X(j-1,i)$
$-\left\{\frac{2j+K+1}{j+1}\right\}\sum_{i=0}^{\infty}\lambda^{i+1}X(j+K,i)$
It can be checked that the $\lambda$ independent terms in eqn.(202) reproduce eqns.(192-93) as required and these correspond to observables for the SHO.
For, non-zero values of $\lambda$ i.e. for $i$=1,2,3,…. one gets from (eqn.202), the following recursion relation for $X(j+1,i)$ by the following procedure: Rewrite (202) by changing $i\rightarrow(i-1)$ in the last-term) leads to
$\sum_{i=1}^{\infty}\lambda^{i}\{X(j+1,i)-a(j)\sum_{m=0}^{i}E_{m}X(j,i-m)$
$-b(j)X(j-1,i)+c(j)X(j+K,i-1)\}$ $=$ $0,~{}~{}~{}~{}~{}~{}$ (205)
Then equating coefficient of each power of $\lambda$ to zero, in (203) one arrives at:
$X(j+1,i)~{}=~{}a(j)\sum_{m=0}^{i}E_{m}X(j,i-m)$
$+b(j)X(j-1,i)-c(j)X(j+K,i-1),$ (206)
where we have denoted:
$a(j)$ $=$ $\left(\frac{2j+1}{j+1}\right),$
$b(j)$ $=$ $\left(\frac{j(4j^{2}-1)}{4(j+1)}\right),$
$c(j)$ $=$ $\left(\frac{2j+K+1}{j+1}\right),$
Finally, it is convenient to change $j\rightarrow(j-1)$ through out in eqn.(205) to get
$X(j,i)$ $=$ $\left\{(\frac{2j-1}{j})\right\}\sum_{m=0}^{i}E_{m}X(j-1,i-m)$
+$+\left\{\frac{(j-1)(4(j-1)^{2}-1)}{4j}\right\}X(j-2,i)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$-\left\{\frac{2j+K-1}{j}\right\}X(j+K-1,i-1).$
Using eqn.(206), along with the result from the FHT given by (see, eqn.(198)), i.e.
$E_{p}~{}=~{}\left(\frac{1}{p}\right)X(K,p-1),~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}{(198)}$ (209)
one can completely solve the above recursion relations for $E_{p}$; $p$=1,2,3,….; provided, the following boundary conditions are imposed,
$X(0,i)~{}=~{}\delta_{0,i}~{}~{}~{}~{},~{}~{}~{}~{}~{}X(j,i)~{}=~{}0~{}for~{}i<0.$ (210)
which follow trivially from the definitions of $X(0)=1$ and $i>0$ respectively.
In respect of the recursion relation, eqn.206 the following observations are in order:
(a) it may be noted that the recursion terminates only when second argument of the last-term in eqn.(206) becomes zero due to the boundary conditions, eqn.(207). This implies that for evaluation of $E_{p}$, one needs to compute all coefficients $X(j,i)$ for $i=1,2,3,....(p-1)$; and for each fixed $i~{},~{}~{}j=1,2,3,......{(p-i)(K-1)+1}$ such that the “maximum” value for ‘${}j~{}$’ goes with “minimum” value of ‘${}i~{}$’. This specifies the set of values of ${}i,j~{}$ over which the recursion in eqn.(206) is to be carried out.
(b) The array of $X(j,i)$ required to solve the problem for computation of $E_{p}$ (i.e.$p$-th perturbation correction) can be generated by both numerical or symbolic implementation of computation. However, it soon becomes unwieldy in ’numerical’ implementation when $p\gg 1$ , due to large rounding-off errors and memory-restrictions. In ref.[9],symbolic implementation through “MAPLE” has been advocated, which we have followed and checked its suitability for $p\leq 200$, which suffices for most theoretical purposes (such as studies of large order behavior of SFPT).
(c) Sample results in SFPT for the QAHO were obtained by the recursion method, eqn.(206), which correctly reproduces [9] the results from WKB-asymptolic computation of Bender-Wu, Banks,Bender and Wu [88,100] for establishing the large-order-behavior confirming the Asymptotic nature and the predicted divergence in SFPT.
(d) More importantly, the“Hypervirial-Feynman-Hellman-Theorem-(HVT-FHT)” - Method, followed here, provides exact (i.e. no truncation) results at each order of perturbation theory.
XIV Application of MFPT to anharmonic-interactions
In the previous section, the method based upon the Hypervirial- and the Feynman-Hellman Theorems (HVT-FHT) for computing perturbation corrections in the “standard formulation of perturbation theory”(SFPT) was highlighted and its supremacy over the conventional method was demonstrated. The ease and simplicity of implementation of the same was established by reproduction of the results for perturbation corrections of arbitrary order in SFPT for the different cases of anharmonic oscillator obtained by other approaches.
In the following sub-sections, we carry out the application of the HVT-FHT method in the framework of the “Mean-Field Approximation Scheme” and the “Mean Field Perturbation Theory” (MFPT) to the case of the QAHO, SAHO and QDWO, which is among the main objectives of this work , in view of the inadequacies of SFPT stated earlier ( see, Sections -II, IV and XI).
XIV.1 The Quartic- and Sextic Anharmonic Oscillators (AHO)
For both these cases,the system-Hamiltonian is given by:
$H~{}=~{}\frac{1}{2}p^{2}+\frac{1}{2}x^{2}+gx^{2K},\mbox{(}K=2,3),$ (211)
corresponding to the quartic-AHO (QAHO) and the sextic-AHO (SAHO) respectively.To carry out the NGAS, we choose [82] the harmonic-approximation for the “Approximating Hamiltonian” (AH), given by eqn.(25 but reproduced below for ready reference:
[…]
which is the Hamiltonian of an energy-shifted-harmonic-oscillator involving two free-parameters:$h_{0}$ and $\omega$, corresponding to an ‘energy-shift’ and the ‘frequency’ respectively. These parameters are determined [82] in accordance to the procedure through the steps(i)-(iii)as outlined in Section-VIA above. For the case of the QAHO, these are determined by equations(29-30) but reproduced here below for ready reference:
[…]
and
[…]
where, the notation is : $\xi\equiv(n+\frac{1}{2});n=0,1,2,3,...$, being the spectral-level index. As derived earlier (in Section-VI-A), these parameters acquire the required functional dependence on $g$ and $\xi$:${}\omega$=$\omega(g,\xi$); ${}h_{0}=h_{0}(g,\xi)$. Similarly, the eigenvalue of the MFH, $E_{0}$ is obtained as a function of $g$ and $\xi$ and given by eqn.(32) (reproduced here, for ready reference)
[…]
(Let us note that because of the imposed constraint of CEQA ( see, Section-VI, eqn.(16)), one can also designate $H_{0}$ as the ”Mean Field Hamiltonian (MFH)”). In Section-VIIB, it has been already discussed that the leading-order(LO) result given by eqn.(32), provides [82] constitutes the dominant contribution to the energy-spectra over of the QAHO over the entire physical range of $n$ and $g$. Besides, it also reproduces the rigorous results [11] on analytic-properties of $E(g,\xi)$ in the complex-$g$ plane.
The next task is to develop the ”mean field perturbation theory (MFPT)” for the QAHO by defining the” perturbation-Hamiltonian as given generally by eqn.(21). For the case of the QAHO. this yields:
$H^{\prime}~{}\equiv~{}H-H_{0}=gx^{2K}-\frac{1}{2}(\omega^{2}-1)x^{2}-h_{0}$ (212)
To determine the perturbation-corrections to the LO-energy-spectrum, we use the ”Associated Hamiltonian (ASH)” introduced earlier (See, eqn.(162)), which now reads as:
$\bar{H}=\frac{1}{2}p^{2}+\frac{1}{2}\omega^{2}x^{2}+h_{0}+\eta(gx^{2K}-\frac{1}{2}(\omega^{2}-1)x^{2}-h_{0})$
$\equiv~{}H_{0}+\eta~{}H^{\prime}.$ (213)
The application of HVT to $\bar{H}$ proceeds in analogous manner as outlined in the previous Section and leads to the following equation :
$2(2j+1)\bar{E}X(j)-2(j+1)\omega^{2}X(j+1)$
$-2(2j+1)h_{0}X(j)+\frac{1}{2}j(4j^{2}-1)X(j-1)$
$+\eta[2(j+1)\alpha X(j+1)+2(2j+1)h_{0}X(j)$
$-2g(2j+K+1)X(j+2)]=0$ (214)
In the eqn.(211) above, we have used the following notation: $K$ is the anharmonicity-index,
$X(j)\equiv\langle\bar{\psi}|x^{2j}|\bar{\psi}\rangle;$ (215)
$\alpha\equiv(\omega^{2}-1)$ and $\bar{\psi}$ , $\bar{E}$ have been defined earlier (see, eqn.(163) ). Next, using the power-series expansion in $\eta$ for the $\eta$-dependent quantities in eqn.(211) and equating the coefficients of each power of $\eta$ to zero,one arrives at the following recursion-relation:
$X(j,i)=a(j)X(j-1,i)+b(j)\sum_{m=0}^{i}E_{m}X(j-1,i-m)$
$+c(j)X(j-2,i)+d(j)X(j-1,i-1)$
$+e_{\omega}X(j,i-1)-f(j)X(j+K-1,i-1),$
where, $X(j,i)$ are defined by the relation: $X(j)=\sum\limits_{i=0}^{\infty}{\eta^{i}}X(j,i)$ and other quantities are given as follows:
$~{}~{}~{}~{}a(j)=\left(\frac{-h_{0}}{\omega^{2}}\right)\left(\frac{2j-1}{j}\right),$ (217a)
$b(j)=\left(\frac{2j-1}{\omega^{2}j}\right),~{}~{}~{}~{}$ (217b)
$c(j)=\left(\frac{(j-1)\left(4(j-1)^{2}-1\right)}{4j\omega^{2}}\right),$ (217c)
$d(j)=\left(\frac{h_{0}}{\omega^{2}}\right)\left(\frac{2j-1}{j}\right)=-a(j),$ (217d)
$e_{\omega}=\frac{(\omega^{2}-1)}{\omega^{2}},$ (217e)
$~{}~{}~{}f(j)=\left(\frac{g}{\omega^{2}}\right)\left(\frac{2j-1+K}{j}\right).$ (217f)
Similarly, following the analogous procedure, as in the previous Section, the Feynman-Hellman Theorem (FHT) provides the following additional relations between $X(j,i)$ and $E_{p}$:
$E_{1}=gX(K,0)-\frac{1}{2}(\omega^{2}-1)X(1,0)-h_{0}~{},~{}~{}~{}~{}$ (218a)
$pE_{p}=gX(K,p-1)-\frac{1}{2}(\omega^{2}-1)X(1,p-1);p=2,3,4,...$ (218b)
[…]
For the case of the QAHO, we note the following results for $E_{p}(g,\xi)$ at $g=1$ and $\xi=\frac{1}{2}$ (i.e. for the ground-state), for some low-lying values of $p$: $E_{1}=0$, $E_{2}=-\frac{3}{256}$,$E_{3}=\frac{27}{4096}$, $E_{4}=-\frac{2373}{262144}$, $E_{5}=\frac{65457}{4194304}....$ etc. (Note that consistent with eqn.(20), we recover the result :$E_{1}=0$ in the limit: $\eta~{}\rightarrow~{}1~{}$, which is of special significance- see later. )
Perturbation corrections for the Sextic-Anharmonic Oscillator (SAHO)
Analogous results for the SAHO can be obtained by simply substituting $K=3$ in eqs.(213-215) and using the following input [82] for $\omega$, $h_{0}$ and $E_{0}$: $\omega$ is given by the real, positive root of the equation, (see, Section-VIID)
$\omega^{4}-\omega^{2}-(\frac{15}{4})(5+4\xi^{2})=0;$ (219)
$E_{0}=(\frac{\xi}{3})(2\omega+\frac{1}{\omega}),~{}~{}~{}h_{0}=(\frac{\xi}{3})(\frac{1}{\omega}-\omega).$ (220)
Some sample-values of $E_{p}$ for the ground state i.e.$\xi=\frac{1}{2}$ and for $\omega=2$ are the following: $E_{1}=0$, $E_{2}=-\frac{49}{960}$, $E_{3}=\frac{671}{4608}$, $E_{4}=-\frac{53621891}{55296000}$, $E_{5}=\frac{2610955409}{265420800}....$ etc. As expected [100,102], the resultant asymptotic series exhibits more severe divergence at large orders as compared to that of the QAHO.
In the next sub-section, we discuss the application of MFPT to the case of the quartic-double-well oscillator (QDWO).
XIV.2 The Quartic Double Well Oscillator (QDWO)
The QDWO constitutes perhaps the simplest case displaying degeneracy of the ground-state. As such, it has played a very important role in the elucidation [82,95,98] of the properties of standard formulation of perturbation theory (SFPT) . In particular,as has been mentioned earlier, successive terms in the perturbation series (PS) of SFPT for the energy eigenvalue are found to be non-alternating in sign at large orders, which frustrates Borel-summability [82,95,98].Several modern developments such as the theory of resurgence [95-99], trans-series [97,98] etc are basically motivated to surmount such typical problems of Borel-nonsummability arising in SFPT when dealing with systems having degenerate ground states.In view of the above scenario,we consider it specially relevant to investigate the QDWO-problem in the frame-work of MFPT.
The Hamiltonian in this case, is given by (see, Section-VIIC):
[…]
The“Mean-Field Hamiltonian” (MFH) is chosen in this case as:
[…]
This choice is motivated to account for the spontaneous symmetry breaking (SSB) through a non-zero vacuum-expectation-value (VEV) of the field operator $\hat{x}$, denoted here by $\sigma$. The parameters $\omega$, $\sigma$ and $h_{0}$ are determined in analogous manner as in the case of the QAHO. The LO-results [82] (See, Section-VIIC) imply ”quantum-phase transition’(QPT) characterized by a critical-coupling:$g_{c}(\xi)$ given by the expression:
$g_{c}(\xi)=\frac{(2/3)^{3/2}}{3(5\xi-(1/4\xi))}$
such that the ‘SSB-phase’ is realized with $\sigma\neq 0$ for $g\leq g_{c}(\xi)$, whereas the ‘symmetry-restored(SR)-phase’ is obtained with $\sigma=0$ when $g>g_{c}(\xi)$. The transition across $g=g_{c}(\xi)$ being discontinuous, the two phases are governed by distinct expressions for $\omega$ and $E_{0}$ ,which are not connected analytically . We, therefore, consider the two-phases separately in the following:
The SR-Phase of the QDWO
In this phase (See, Section-VIIC) we have: $g>g_{c}(\xi)$; $\sigma=0$; $\omega$ satisfies the equation,
[…]
The energy eigenvalues and the energy-shift parameter are given in LO as:
[…]
and  $h_{0}=E_{0}-\omega\xi$
It has been demonstrated in Section-VII and ref. [82] that LO-results for $E_{0}$ as given above, provide accurate estimate of the energy-spectrum to within a few-percent of the ‘exact’-result for arbitrary value of $g$ and $\xi$, thus constituting the major contribution to it. Further improvement over the LO-result can be attempted by the application of MFPT following the procedure analogous to that for the QAHO. For this purpose, the ’perturbation-Hamiltonian’ and the ’associated- Hamiltonian’ are now respectively given by:
$H^{\prime}~{}=gx^{4}-\frac{1}{2}(\omega^{2}+1)x^{2}-h_{0},$ (221)
$\bar{H}=\frac{1}{2}p^{2}+\frac{1}{2}\omega^{2}x^{2}+h_{0}+\eta(gx^{4}-\frac{1}{2}(\omega^{2}+1)x^{2}-h_{0}).$ (222)
[…]
${}E_{1}=gX(2,0)-\frac{1}{2}(\omega^{2}+1)X(1,0)-h_{0}~{},$ (223)
${}pE_{p}=gX(2,p-1)-\frac{1}{2}(\omega^{2}+1)X(1,p-1);~{}~{}p=2,3,4,....$ (224)
[…]
We next consider the SSB-phase of the QDWO.
The SSB-Phase of the QDWO
As stated earlier (Section-VIIC), a quantum- phase-transition(QPT) to this phase occurs when $g<g_{c}(\xi)$ manifesting in degenerate global-minima of the effective potential located at $x=\pm\sigma$, where ${}~{}~{}\sigma^{2}=(1-{12g{\xi}}/{\omega})/4g.$ In this case, $\omega$ satisfies[82] the equation,
${}~{}~{}~{}\omega^{3}-2\omega+6g(5\xi-\frac{1}{4\xi})=0.$ (225)
The LO-result for the energy-spectrum is given by [82]
${}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}E_{0}={(-1/16g)}+{(\xi/4)}{(3\omega+(2/\omega))}~{}$
and the energy-shift is given by: ${}h_{0}=E_{0}-\omega\xi$. Again, the LO-result for $E_{0}$ as given above, provides [82] accurate estimate of the energy-spectrum to within a few-percent of the ‘exact’-numerical result for arbitrary values of $g$ and $\xi$. Further improvement in accuracy can be achieved by computing perturbation-corrections to the LO-result via MFPT as described below: The perturbation-Hamiltonian and the auxiliary Hamiltonian (AH) are respectively given by,
$H^{\prime}~{}=gx^{4}-\frac{1}{2}(\omega^{2}+1)x^{2}+\omega^{2}x-\left(h_{0}+\frac{1}{2}\omega^{2}\sigma^{2}\right),$ (226)
$\bar{H}=\frac{1}{2}p^{2}+\frac{1}{2}\omega^{2}\left(x-\sigma\right)^{2}+h_{0}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$+\eta(gx^{4}-\frac{1}{2}(\omega^{2}+1)x^{2}+\omega^{2}\sigma x-h_{0}-\frac{1}{2}\omega^{2}\sigma^{2}).$ (227)
Due to the presence of odd-powers of $x$ in the expression of $\bar{H}$ as given above, it is appropriate to choose:$f(x)=x^{j},j=0,1,2,3,..$ for the application of HVT to $\bar{H}$. This leads to the following recursion relation [110,111]:
$X(j,i)=a(j)X(j-1,i)+b(j)\sum_{m=0}^{i}E_{m}X(j-2,i-m)$
$-\tilde{b}X(j-2,i)+c(j)X(j-4,i)+d(j)X(j-1,i-1)$
$+\tilde{b}X(j-2,i-1)+e_{\omega}X(j,i-1)-f(j)X(j+2,i-1),$
where, $X(j)\equiv\langle\bar{\psi}|x^{j}|\bar{\psi}\rangle$,${}X(j,i)$ are defined by the relation: $X(j)=\sum\limits_{i=0}^{\infty}{\eta^{i}}X(j,i)$
and other quantities are given as follows:
$a(j)=\sigma\left(\frac{2j-1}{j}\right),~{}~{}~{}~{}~{}~{}~{}~{}$ (229a)
$b(j)=2\left(\frac{j-1}{\omega^{2}j}\right),~{}~{}~{}~{}~{}~{}~{}~{}$ (229b)
$\tilde{b}(j)=(h_{0}+\frac{1}{2}\omega^{2}\sigma^{2})b(j),~{}~{}~{}~{}~{}~{}~{}~{}$ (229c)
$c(j)=\frac{(j-1)(j-2)(j-3)}{\left({4j\omega^{2}}\right)},~{}~{}~{}~{}~{}~{}~{}$ (229d)
$d(j)=-a(j),~{}~{}~{}~{}~{}~{}~{}$ (229e)
$e_{\omega}=\frac{(\omega^{2}+1)}{\omega^{2}},~{}~{}~{}~{}~{}~{}~{}$ (229f)
$f(j)=\left(\frac{2g}{\omega^{2}}\right)\left(\frac{j-1+K}{j}\right).~{}~{}~{}~{}~{}~{}$ (229g)
Application of FHT then relates $E_{p}$ to the $X(j,i)$ as follows:
$E_{1}~{}=~{}gX(4,0)-\frac{1}{2}(\omega^{2}+1)X(2,0)+$
$\omega^{2}\sigma X(1,0)-\left(h_{0}+\frac{1}{2}(\omega^{2}\sigma^{2})\right),$ (230)
and
$pE_{p}~{}=~{}gX(4,p-1)-\frac{1}{2}(\omega^{2}+1)X(2,p-1)$
$+(\omega^{2}\sigma)X(1,p-1),~{}~{}$ (231)
where $p=2,3,4....$
Note that the computation of $E_{p}$ using the eqns.(225-228) above, requires that the recursion in eqn.(225) traverses within the following range of values of $i,j$ (in steps of unity):
$0\leq~{}i~{}\leq(p-1),~{}and~{}1\leq~{}j~{}\leq 2(p+1-i).$ (232)
With the above inputs and boundary conditions, the perturbation corrections $E_{p}$ can, in principle, be computed exactly to arbitrary high order $p$. We note the following results [110,111] for $E_{p}$ at $\xi=\frac{1}{2}$ and $\omega=1$ , which implies $g=(1/12)<g_{c}$ as required:
$E_{1}=0$, $E_{2}=-\frac{17}{384}$, $E_{3}=\frac{83}{3072}$, $E_{4}=-\frac{69943}{884736}$, $E_{5}=\frac{464195}{2359296}....$ etc.
[…]
XV Computation of the Total Perturbation Correction ( TPC )
We compute the TPC by two main methods: (a) the method of ‘optimal truncation’ (MOT) of the original asymptotic perturbation series and (b) by ’Borel-summation’. These are sequentially described in the following sections. (For convenience and illustration, we confine here to the computation for the ground state.)
XV.1 Method of Optimal Truncation (MOT)
This method utilizes the property [108], that the initial terms of an asymptotic series(AS) continue to decrease in magnitude till the “term with the least magnitude (TLM)” is reached, beyond which, the subsequent terms exhibit monotonic increase. Hence, the AS can be reasonably approximated by truncation at the TLM.
In the case of “Mean Field Perturbation Theory (MFPT)”, in view of eqn.(161), we have:
$\left(E(g)-\sum\limits_{k=0}^{N}{E}_{k}(g)\right)\leq{{E}_{N+1}(g)}.$ (233)
Therefore, if $N_{0}(g)$ denotes the TLM then the MOT leads to the following estimate:
${}~{}E(g)~{}\simeq\sum\limits_{k=0}^{N_{0}(g)}{E}_{k}(g).$ (234)
This method (MOT) works for arbitrary physical value of $g$ in MFPT. In Table-I, we present the results from this method for the case of the QAHO, QDWO and SAHO for sample-values of $g$. It can be seen from this Table that the TLM occurs at reasonably low values : $N_{0}(g)\sim$ (2-6).The primary reason for this occurrence can be traced back to the fact that $E_{1}(g)$ vanishes for arbitrary $g$ and since the terms of the ’asymptotic perturbation series’ must diverge at large orders, the TLM is constrained to occur at fairly low-values, without compromising the accuracy , as exhibited in Table-I.
In this Table-I, we have listed computed values for the ’total perturbation correction (TPC) ’ by the method of ’optimal truncation(MOT)’ of perturbation series(APS) in MFPT, for the ground-state energy at different values of the coupling-strength $g$ (col.(1)). Results are shown (col.(4)) for the ground-state energy of QAHO, SAHO and QDWO and compared with accurate numerical results of earlier calculations denoted as `$Exact^{\prime}$ (col.(5)), from the given reference. The LO-results in MFPT are given in col.(2). In col.3, $N_{0}(g)$ denotes the ”truncation-point” of the respective APS at the‘term of least magnitude’(TLM). Accuracy as percentage-errors are depicted in the last-column relative to the‘exact’-results. For comparison with ‘standard’ results, we have considered earlier computations in [112],[113] and [114] for the case of QAHO, SAHO and QDWO respectively. The relative-error incurred is seen to be within a few-percent uniformly over the considered range of $g$.
In view of these results, the MOT in MFPT may be regarded as a ‘fail-safe’ method in achieving reasonable accuracy for arbitrary values of $g$. This is because the MOT remains as the only practical method available when other methods , such as Borel-summation etc fail or are inapplicable.[115]
TABLE- I
$QAHO$ $ref.(112)$ $g$ $E_{0}$ $N_{0}(g)$ $E_{MOT}$ $Exact$ $Error{(\%)}$ 0.1 0.5603 6 0.5593 0.5591 0.03 1.0 0.8125 3 0.8074 0.8038 0.44 10.0 1.5312 3 1.5204 1.5050 1.02 100.0 3.1924 3 3.1701 3.1314 1.23 $SAHO$ $ref.(113)$ 0.1 0.5964 2 0.5787 0.5869 1.40 1.0 0.8378 2 0.7694 0.8050 4.42 50.0 1.9735 2 1.7241 1.8585 7.23 200.0 2.7606 2 2.3986 2.5942 7.54 $QDWO$ $ref.(114)$ 0.1 0.5496 2 0.4107 0.4709 12.78 1.0 0.5989 3 0.5998 0.5773 3.91 10.0 1.4097 3 1.4007 1.3778 1.66 100.0 3.1338 3 3.1122 3.0701 1.37
TABLE-I : See text below eqn.(231) for a description of entries in this Table.
XV.2 Method of Borel Summation (MBS)
Consider the generic case ( see, eqn.(161)) when we have $E_{j}\sim~{}\Gamma(\alpha j+1)$ for $j\gg~{}1$. Then, it follows that
$\lim_{j\gg 1}b_{j}\equiv\dfrac{E_{j}}{\Gamma(\alpha j+1)}\rightarrow 0,~{}~{}\alpha>0.$ (235)
Using the integral representation of the Gamma function:
[…]
one can formally express the “Total Perturbation Correction” (TPC) $\Delta{E(g)}$ (see, eqn.(161)) as :
[…]
where, we have defined,${}u=t^{\alpha},\gamma=\frac{1}{\alpha}$ and $B(u)$ denotes the ‘Borel-Series(BS)’, given by :
$B(u)\equiv\sum\limits_{j=1}^{\infty}b_{j}u^{j}.$ (238)
Note that, although by virtue of eqn.(232), $B(u)$ has a finite ‘radius of convergence’ $r_{c}$ , the ‘Borel-Laplace-integral’(BLI), eqn.(233) does not exist yet since the range of integration extends beyond $r_{c}$.
The way out of this difficulty is to carry-out the analytic continuation/ extrapolation of the BS to the full-path of integration of the BLI, i.e. to the whole $+ve$-real axis in the u-plane. Denoting such analytic-continuation by $\tilde{B}(u)$ and substituting this for $B(u)$ in eqn.(234), the BLI can now be made to exist and hence can be used to‘define’ the TPC as:
[…]
The feasibility of Borel-summation thus crucially depends upon the success of the analytic continuation/ extrapolation with the afore stated properties. In practice however, $\tilde{B}(u)$ can only be inferred approximately from the BS , if feasible, since only a finite number of terms in eqn.(235) are computable/known. Under such circumstances, the method of conformal mapping (MCM) has been preferably used [116] for implementing analytic continuation in a model-independent way. We, therefore employ the same in the following.
The inputs, which are required for the implementation of MCM are the location and nature of the singularity lying closest to the origin, which we designate as the ‘leading singularity’(LS). By the ‘Darboux theorem’ [117], the late-terms in the BS, eqn.(235), originate from the LS. Therefore, the nature and location of the LS may be inferred from the terms of the BS at large orders. This is readily demonstrated considering the following ansatz of the LS: $\tilde{B}(u)\simeq\tilde{B}_{0}(u)\sim~{}(u+r_{c})^{p}$. The ‘radius of convergence’ $r_{c}$ and the ‘singularity-exponent’ $p$ can then be determined from a fast-converging set of equations [9] given below:
${}~{}~{}~{}r_{c}(g)=\lim_{j\gg 1}~{}\dfrac{b_{j}b_{j-1}}{jb_{j}^{2}-(j+1)b_{j+1}b_{j-1}},$ (240)
$p(g)=\lim_{j\gg 1}~{}\dfrac{{jb_{j}^{2}}-(j^{2}-1)b_{j-1}b_{j+1}}{jb_{j}^{2}-(j+1)b_{j+1}b_{j-1}}.$ (241)
We obtain the value:${}p(g)=~{}-0.5$, being independent of $g$ as expected and the values for $r_{c}(g)$, as shown in Table-II. Equipped with these information as inputs, consider [118] the conformal mapping , $z$ which maps the cut-u plan,$|arg(u)|<\pi$ into the interior of the unit-disk in the $z$-plane while preserving the origin:
$z(u)=\dfrac{\sqrt{1+su}-1}{\sqrt{1+su}+1};~{}~{}s=\dfrac{1}{r_{c}};~{}~{}u=t^{\alpha},$ (242a)
and its inverse
$u(z)=\left(\frac{4}{s}\right)\dfrac{z}{(1-z)^{2}};~{}~{}|z|\leq 1.~{}~{}~{}~{}$ (242b)
Note that the images of the points: $u=\pm\infty$ are at $z=1$ and further that $z(u)$ is a real-analytic-function of u.
The other important property of the conformal-transformation, eqn.(239) is its ‘optimality’ as established by the ‘Ciulli-Fischer Theorem’ [118]. According to this theorem, the error of truncation becomes minimum when an arbitrary function $f(z)$ expanded as an infinite- power-series in $z$, is approximated by truncation at a finite number of terms, thus ensuring the optimal rate of convergence.
Hence, consider the partial sum of the BS, eqn.(235):  $B_{N}(u)\equiv\sum\limits_{k=1}^{N}b_{k}u^{k}$.
Using the inverse-transformation, eqn.(239b),$B_{N}(u)$ can be consistently re-expanded [119,120] as follows, noting that $u^{N}\sim~{}z^{N}+O(z^{N+1})$:
$B_{N}(u)\equiv\sum\limits_{k=1}^{N}b_{k}u^{k}\rightarrow~{}\bar{B}_{N}(z)=\sum\limits_{k=1}^{N}B_{k}z^{k},$ (243a)
[…]
where, we have denoted the analytic-continuation of the partial sum of the BS by:${}B_{N}(u)\rightarrow~{}\bar{B}_{N}(z)$.
Using the above properties/results, the TPC can be evaluated by the change of the integration-variable to $z$ (see, eq(239b)).Then, by taking care of possible numerical-divergence at the upper limit, ${}z=1$ and further, considering the limit of the sequence of partial-sums, the TPC is evaluated as follows ( we suppress the $g$-dependence of quantities for notational-clarity):
[…]
 with $\mbox{~{}with~{}}f_{N}(z)=(\gamma\rho)~{}\dfrac{(1+z)}{(1-z)^{3}}\left(\dfrac{\rho~{}z}{(1-z)^{2}}\right)^{\gamma-1}$
$~{}exp\left[-\left(\dfrac{\rho~{}z}{(1-z)^{2}}\right)^{\gamma}~{}\right]\bar{B}_{N}(z),$ (244b)
where,${}\rho=4r_{c}$ and ${}\bar{B}_{N}(z)$ is defined by eqn.(240a). Note that term-by-term integration of the partial-sum,${}\bar{B}_{N}(z)$ can now be carried out in eqn.(241a), in contrast to the BS in eqn.(234) because of the incorporation of analyticity through the conformal-map $z$.
TABLE II
$g$ $r_{c}$ $N_{c}$ $\varDelta{E(g)}$ $E_{0}$ $E_{tot}(g)$ $Exact$ $Er({\%})$
QAHO $(\alpha=1)$ $ref.(112)$
0.1 6.071 6 -0.00116 0.5603 0.55914 0.5591 0.0083
1.0 2.667 7 -0.00869 0.8125 0.80381 0.8038 0.0003
10.0 2.133 8 -0.02619 1.5312 1.50501 1.5050 0.0005
100.0 2.028 10 -0.06101 3.1924 1.13139 3.1314 0.0001
SAHO $(\alpha=2)$ $ref.(113)$
0.1 13.3 20 -0.0095 0.5964 0.5869 0.5869 0.001
1.0 8.56 20 -0.0328 0.8378 0.8050 0.8050 0.002
10.0 7.14 20 -0.1149 1.9735 1.8586 1.8585 0.007
100.0 7.02 20 -0.1662 2.7606 2.5944 2.5942 0.007
QDWO $(\alpha=1)$ $ref.(114)$
0.5 1.191 20 -0.0232 0.4770 0.4538 0.4538 0.0027
1.0 1.455 11 -0.0216 0.5989 0.5773 0.5773 0.0006
10.0 1.872 20 -0.0320 1.4098 1.3778 1.3778 0.0040
100.0 1.972 18 -0.0637 3.1338 3.0701 3.0701 0.0005
TABLE-II: See text following eqn.(241b), for a description of the entries in this Table
The results are presented in Table-II. There we have shown ( col.(4)) the Borel-sum for the total- perturbation-correction(TPC) computed using eqn.(241a-b) and conformal-mapping, eqn.(239a) , for the ground state energy of the QAHO, SAHO and the QDWO as a function of $g$ given in col.(1). For each case, the input-value of $\alpha$, (see eqn.(232)) is indicated. Entries in col.(2) are the input-value for the ‘radius of convergence’; $N_{c}$ in col.(3) denotes the number of terms retained in the Borel-series for convergence up to the specified accuracy; $E_{0}$ in col.(5) denotes the LO-contribution; $E_{tot}$ in col.(6) represents the ‘total corrected energy’; entries in col(7) labeled $Exact$, display the ‘standard’-numerical results from the indicated reference; and the last column specifies the %${\%}error$ with respect to the $Exact$-result. We have fixed the value:  $\epsilon=0.001$ ,( see eqn.(241a)) in the computation of TPC in all-cases.
It may be noted that the convergence of the partial-sum is quite fast- an accuracy of $\sim(99.99\%)$ is uniformly achieved over the full explored range of $g$ and for all the cases of anharmonic-interaction since the cut-off $N_{c}$ on the number of terms retained in the partial-sum is as low as $\sim\left(10-20\right)$ (Table-II). Thus, the ‘$optimality$’ and the ‘$accuracy$’ of the ’method of conformal mapping’ are firmly established in MFPT.
XV.3 Section - Summary
We have presented a new formulation of perturbation theory based upon a mean-field- Hamiltonian-approximation and have demonstrated that it is Borel-summable for arbitrary coupling-strength in the case of QAHO, SAHO and the QDWO, by obtaining very accurate results for the energy-spectrum. By this demonstration, the distinction between the‘perturbative’ and the ‘non-perturbative’-regimes appears to be an artifact based on SFPT.
The results obtained for the QDWO may be of particular significance because this case typically belongs to systems with ‘degenerate ground states’, which are generically found to frustrate Borel-summability in the SFPT. It may also be significant that, results of comparable accuracy have also been obtained [104]( see, Section-VIII ) with a much cruder approximation using the square-well-potential in the ’mean-field Hamiltonian MFH’ which preserves the general features of the ’mean field approximation scheme (MFAS) ’, thus demonstrating the general nature and flexibility of the input in MFAS.
In view of the simplicity and the general-nature of the formulation demonstrated in this investigation , we are inclined to conjecture that MFPT may be applicable to arbitrary interacting systems in quantum-theory.
XVI Summary and Conclusions
In summary, a non-perturbative scheme of approximation in quantum theory, designated herein as: “ Non-perturbative General Approximation Scheme” (NGAS) is presented in this report. This scheme is in principle, applicable to general interacting systems described by a Hamiltonian and has the distinction of being simple in formulation, non-perturbative, self-consistent, flexible to incorporate different choice of approximating input and more importantly, admits of systematic improvement via a perturbation theory, which is valid for arbitrary strength of interaction.
The crucial feature of this scheme consists of finding a “mapping” which maps the “interacting system” on to an “exactly solvable” input-model, while preserving the major effects of interaction through a self consistent feed-back mechanism of demanding equal quantum averages of observables in the two systems.
This approximation method has the advantage over the standard formulation of perturbation theory (SFPT) and the variational approximation by transcending the limitations of both: unlike the variational method, it is systematically improvable through the development of a new perturbation theory namely, the “Mean Field Perturbation Theory”(MFPT), whereas in contrast to the case of the SFPT, MFPT preserves the analyticity and non-linearity of the original system order-by-order, while being Borel-summable, for all allowed values of the coupling strength of interaction. Moreover, better behavior of the divergent asymptotic perturbation series at large orders, is seen to emerge in MFPT.
The scheme reproduces the results obtained by several earlier methods, e.g.[112-114] while transcending the limitations of these methods in respect of wider applicability, systematic improvement and better convergence.
It may also be significant that NGAS allows flexibility in the choice of the approximating Hamiltonian (AH). Although the results presented here are based upon the “harmonic-approximation” , the same with comparable accuracy have also been obtained via a much cruder approximation using the square-well-Hamiltonian as the input Hamiltonian within the framework of the general features of NGAS.
Similarly, in the domain of quantum field theory, the scheme has been applied to the case of $g\phi^{4}$ theory, with the choice of the approximating Hamiltonian as that for the free-hermitian-scalar field (but with adjustable ‘mass’ and ‘shifted’ field). In this case, the standard results of the Gaussian-approximation are reproduced in the LO including the non-perturbative renormalization of the ‘mass’ and ‘coupling strength’. However, NGAS appears to be more general in scope and insight than the Gaussian approximation since the former provides a ’dynamical explanation’ of the latter as emergent through the mechanism of altered vacuum-structure introduced by the interaction. Besides, the scheme goes beyond the scope of the Gaussian-approach in establishing new results, e.g., the calculation of the momentum-distribution of the condensate-structure function ‘$n(k)$’. It may also be emphasized that by going beyond the LO of NGAS, the results of the Gaussian approximation can be systematically improved, order-by-order.
The resulting momentum distribution of the vacuum condensate structure function ‘$n(k)$’ deserves special mention as it displays the non-standard feature of an appreciable spread in ${}|k|~{}$ about the origin, scaled by the renormalized mass of the physical quanta. It is reasonable to expect that this condensate-structure of the physical vacuum persists to exist at finite temperature manifesting in the thermodynamic properties of the system with testable consequences.
As has been noted above, a remarkable feature of NGAS is that, while the LO-results are essentially non-perturbative, yet the same are improvable order-by-order in MFPT. The other features of MFPT, which are distinct from those of the conventional perturbation method,e.g. the former is free from power-series-expansion in any physical parameter, including the coupling strength and applicable for arbitrary strength of interaction.Besides, it has been demonstrated that MFPT is Borel-summable (for arbitrary coupling-strength) in the case of QAHO, SAHO and the QDWO, by obtaining very accurate results for the energy-spectrum.
By circumventing the limitations of SFPT in MFPT, it is demonstrated that the distinction between the‘perturbative’ and the ‘non-perturbative’-regimes vanishes and, therefore the same could be a mere artifact based inthe former formulation i.e. the standard formulation of the perturbation theory (SFPT).
The demonstration of Bore-summability of MFPT for the QDWO may be of special significance and contrasted with those derived in SFPT, because this case typically belongs to systems with ‘degenerate ground states’ and such cases are generically found to frustrate Borel-summability in the SFPT. As discussed in this report, this situation , which has given rise to considerable efforts in recent literature [95-99] to surmount this problem of the SFPT for the QDWO and other such systems , is not warranted in MFPT.
In view of the simplicity and the general-nature of the formulation , it may be reasonable to conjecture that this scheme, NGAS may be applicable to arbitrary interacting systems in quantum-theory.
XVII Outlook
In this report we have confined the application of the scheme to some bench-mark systems in quantum mechanics and quantum field theory, namely the anharmonic interactions in one-dimension including the quartic-, sextic- and octic-anharmonic oscillators , the quartic-double well-oscillators and to the ${}g\phi^{4}~{}$ quantum field theory. The investigation can be extended in different directions, i.e. to include finite temperature field theory, quantum-statistics, non-oscillator systems, super-symmetric cases and quantum field theories involving fermions and gauge-fields etc.
Possible applications of the results derived here are envisaged in diverse area of current interest including critical phenomena (involving a scalar field as the order-parameter [84]), inflationary cosmology [87], finite temperature field theory [121] , exploration of the vacuum structure [91] of pure gluonic-QCD and Higgs sector of the standard model ( by extending [90] the analysis to the spontaneously broken phase, which corresponds to the case of negative bare-mass ${}m^{2}<0~{}$).
One immediate task could be to apply MFPT to other systems known to be Borel non-summable in SFPT [94] to determine whether summability can be restored in such cases. Similarly, tunnel-splitting of energy levels [95] can be studied after computation of perturbation correction to the excitation spectrum of the QDWO using more realistic [122] input - Hamiltonian, if necessary. Application to other Hamiltonian-systems are envisaged in a straight forward manner in view of the universal nature of the approximation scheme.
Acknowledgments
The authors acknowledge useful comments and suggestions from the anonymous referees of the journals at references [48],[50],[82] and[110] . The computation and library facilities at the Institute of Physics, Bhubaneswar (IOPB) and NISER, Bhubaneswar are gratefully acknowledged where one of the authors , (BPM) held positions of guest-scientist and visiting faculty respectively. The authors are thankful to Prof. F.M.Fernandez for help in symbolic-computation using Maple and for helpful comments. Useful comments from Profs. A. M. Srivastava, S. G. Mishra and P. Agrawall during seminar talks on the topic(s) of this investigation are gratefully acknowledged.